### v2

# Elastic properties and thermodynamic anomalies of supersolids

Abstract
We study a supersolid in the context of a Gross-Pitaevskii theory with a non-local effective potential. We employ a homogenisation technique which allows us to calculate the elastic moduli, supersolid fraction and other state variables of the system. Our methodology is verified against numerical simulations of elastic deformations. We can also verify that the long-wavelength Goldstone modes that emerge from this technique agree with Bogoliubov theory. We find a thermodynamic anomaly that the supersolid does not obey the thermodynamic relation $\partial P/\partial V|_{N}\,=-n\,\left(\partial P/\partial N|_{V}\right)$, which we claim is a feature unique to supersolids.
I Introduction
A supersolid is a phase of matter that displays both crystalline order and superfluidity in the form of non-classical rotational inertia (NCRI). A key requirement is that both the continuous translational symmetry and the U(1) global gauge symmetry of the system is spontaneously broken in the ground state.
There have been several attempts to understand this quantum phase of matter experimentally and to understand its properties theoretically. While it has not been observed in bulk 4He [1, 2], there are hints that such a phase may exist in the second monolayer of 4He on graphite [3] where NCRI has been measured in a density regime near layer completion with an anomalous temperature dependence of the specific heat capacity. More recent experiments in ultracold dysprosium and rubidium atoms [4, 5, 6] have observed the spontaneous breaking of continuous translational symmetry together with long-range phase correlation.
The first considerations of a supersolid phase were made by Andreev and Lifshitz [7] and Chester [8] who considered the possibility of a supersolid phase in 4He. They argued that the superfluid fraction of a supersolid would be reduced from 100% due to the coupling of the phonons of the crystalline structure to the $U(1)$ phase of the condensate wavefunction. This was further developed by Leggett [9].
Later attempts at theoretical work have taken a phenomenological symmetry-based approach starting with the work of Nozières [10] and Dorsey et al. [11]. Son [12] has described how Galilean invariance puts constraints on the form of the Lagrangian that describes the low-energy dynamics of supersolids.
There have also been approaches starting from microscopic Hamiltonians based on Gross-Pitaevskii theory [13] and Bogoliubov theory [14]. It is natural to ask whether these approaches give rise to the same predictions for the low-energy properties of the supersolid system, such as elasticity and superfluidity.
Moreover, there has been intensive study on the superfluid fraction and excitation spectrum of the supersolid phase. However, there has been comparatively little study on the elastic properties. In this paper we will build on the homogenisation technique of Rica and co-workers [15, 16, 17] to obtain an effective low-energy theory that agrees well with Bogoliubov theory and numerical calculations. Importantly, the original formulation of the homogenisation technique obtained a value for the bulk modulus that was not in agreement with the expected bulk modulus in the superfluid phase. Furthermore, a triangular supersolid would be expected to possess certain symmetries in the Cauchy elastic tensor, and these were not satisfied in the original formulation. Finally, the velocities of the long-wavelength excitations in the previous work also did not agree with the results based on Bogoliubov theory [14], in the sense that the velocities obtained on a particular supersolid phase using the Bogoliubov technique were not consistent with those obtained via the original formulation of homogenisation.
With some important corrections that will be derived in this paper, we will show that the homogenisation approach can be reconciled with other techniques, producing the predictions for all the elastic and superfluid properties of the supersolid phase. More specifically, we provide a method with the Gross-Pitaevskii approximation that calculates the elastic constants and superfluid density (phase stiffness) which is valid for any particular ground-state solution of a Gross-Pitaevskii Lagrangian. It should be noted that, in the limit of zero temperature, where a homogeneous superfluid has a superfluid fraction of 100%, a supersolid has a reduced superfluid fraction [9, 18]. It is important to clarify that this is not an enhancement of the “normal fraction” which is found in superfluids at non-zero temperatures and carries entropy. Rather, it is indicative of a separate phase that spontaneously breaks translational symmetry. In fact, it is a consequence of the fact that the “phonons” of the crystalline structure couples to the $U(1)$ phase of the condensate wavefunction. [19, 20, 21, 22, 23].
The outline of the paper is as follows. In section 2, we review the Gross-Pitaevskii theory for Bose condensates and how a finite-range interaction can give rise to a supersolid phase. In section 3, we lay the framework for the homogenisation theory, specifying the deformation procedure and rigorously defining all necessary steps to calculate a long-wavelength theory of the low-lying excitations. Section 4 uses results from section 3 to derive elastic constants analytically and finds that they agree with elastic constants that we obtain numerically. In section 5, we will consider the additional $U(1)$ gauge and derive the supersolid fraction, as well as a coarse-grained Lagrangian which is the analytic expansion around the ground-state of the elastic strain and $U(1)$ fields. Section 6 uses the effective Lagrangian and considers the additional energetic contributions of flow across the system, and of work done by the strain on the surroundings. In doing so, we derive a long-wavelength effective Lagrangian which describes the Goldstone modes of the system. Section 7 solves for the excitation velocities and discusses some of the key features of the theory. In section 8, we verify the excitation velocities through use of Bogoliubov fluctuations in a Bloch theorem context, and find excellent agreement between the two seemingly disparate theories. In Section 9, we will discuss a thermodynamic anomaly in the compressibility that we believe is unique to a supersolid.
We subsequently make the claim that since we have satisfied the symmetry requirements for an effective theory imposed by Son [12], have verified our elastic constants numerically, and have verified the velocities of the low-lying excitations through two independent and seemingly unconnected techniques; then we must have the correct effective Lagrangian for a supersolid. Furthermore, the technique outlined in this paper can now be used and applied directly to other more complicated and realistic systems.
II Meanfield Supersolids
In this paper, we study a Bose-Einstein condensate (BEC) of particles of mass $m$ with a finite-range interaction $U(\mathbf{r})$ in a Gross-Pitaevskii theory. The condensate wavefunction $\psi(\mathbf{r},t)$ is a complex-valued function that can be written in number-phase representation (or Madelung form) as $\psi(\mathbf{r},t)=\sqrt{\rho(\mathbf{r},t)}e^{i\phi(\mathbf{r},t)}$. The Lagrangian of the system is
$\begin{split}\mathcal{L}=-\int_{\Omega}&\left[\hbar\rho\frac{\partial\phi}{\partial t}+\frac{\hbar^{2}}{2m}\left(\rho(\nabla\phi)^{2}+\frac{1}{4\rho}(\nabla\rho)^{2}\right)\right.\\ &\left.+\frac{1}{2}\,\rho(\mathbf{r})\int_{\Omega}U(\absolutevalue{\mathbf{r}-\mathbf{r}^{\prime}})\rho(\mathbf{r}^{\prime})\mathop{}\!\mathrm{d}\mathbf{r}^{\prime}\right]\mathop{}\!\mathrm{d}\mathbf{r}\end{split}$ (1)
where $\Omega$ is the spatial domain of the system. In the absence of any current or twisted boundary conditions, the phase $\phi$ is spatially uniform in the ground state and can subsequently be set to zero. The ground state density can be found by variational calculus and is given by
$\frac{\hbar^{2}}{4m}\left(\frac{(\nabla\rho)^{2}}{2\rho^{2}}-\frac{\nabla^{2}\rho}{\rho}\right)+\int_{\Omega}U(\mathbf{r}-\mathbf{r}^{\prime})\rho(\mathbf{r}^{\prime})\mathop{}\!\mathrm{d}\mathbf{r}^{\prime}=\mu\,.$ (2)
where $\mu$ is the chemical potential to enforce the constraint $\int_{\Omega}\rho(\mathbf{r})\mathop{}\!\mathrm{d}\mathbf{r}=N$, the total number of particles in the system Equation 2 is a nonlinear equation which we solve numerically via a Crank-Nicholson scheme. By evolving the Lagrangian in imaginary time numerically, we obtain the (real) ground state wavefunction as the steady-state solution, using a Runge-Kutta algorithm for the temporal evolution. The numerics are performed on a 2D triangular grid in real space with periodic boundary conditions in order to respect the hexagonal symmetry of the spatially modulated condensate that we expect.
A necessary feature (as shown by Heinonen et. al [24]) for the spontaneous development of a density wave in the condensate is the presence of an interaction potential which at least has non-zero Fourier components in $\mathbf{k}^{4}$ and upwards [24]. This rules out the use of the typical contact potential for supersolid formation.
To illustrate this, consider a toy model of a soft-core repulsive interaction with a finite-range of $a$ in two dimensions:
$U(\mathbf{r})=U_{0}\Theta(\absolutevalue{\mathbf{r}}-a)$ (3)
where $\Theta$ is the Heaviside step function. For this simple interaction, the system is controlled by a single dimensionless parameter
$\Lambda=\frac{\pi U_{0}ma^{2}}{\hbar^{2}}na^{2}$ (4)
for a system of average number density $n$.
The phase diagrams for the soft-core system are described in detail in various works [25, 26, 27]. We find that our results are in agreement, with small $\Lambda$ corresponding to a superfluid phase and that for $\Lambda>37$ the system spatially orders into the supersolid phase. This is in agreement with a theoretical subcritical limit calculated by During et al [17]. This supersolid phase is characterised by a triangular lattice in 2D, illustrated in Figure 1.
The bulk of this paper is concerned with an effective long-wavelength theory for this supersolid phase where the Bose condensate is spatially modulated. We will use these numerics to verify our formalism by calculating the elastic constants which are dependent on the ground state.
Figure 1: A spatially modulated condensate. Left: real space density distribution. Right: Fourier transform of density distribution implying non-zero momenta in the condensate.
III Homogenisation for Elasticity
The spontaneously broken translational symmetry in a supersolid means that the system must possess Goldstone modes which are the elastic modes. In this section, we will build on the work of [15, 17] to derive the elastic response of the system. The goal here is to consider an elastic deformation on the ground state, e.g. stretching/shearing as described by some strain tensor, and then to expand the Lagrangian to second order in said strain tensor and recover the elastic moduli through the generalised Hooke’s law.
We follow the metholodology of Josserand et al [15], but our results have some key differences. Namely, we believe some key terms were missed which skews the elastic constants to ones that do not agree with simple calculations in the superfluid limit and lead to physically inconsistent elastic moduli. More specifically, the Cauchy elastic tensor does not obey internal symmetries consistent with a triangular lattice, and possesses negative elastic moduli which suggest an unstable state. Moreover, we conduct a different treatment of the system with regard to the canonical ensemble and the degrees of expansion in the strain tensor.
The analysis will take the following structure: we will first carefully define the notion of an elastic deformation on the system and provide the basis for the calculation. We then expand the Lagrangian to second-order in the elastic strain tensor. We demand that the deformed system also be a ground state, and therefore solve the Euler-Lagrange equations for the new system, which will allow us to determine elastic constants.
We begin by describing a physical deformation by considering the displacement, $u(r)$, of points in the material. The displacement can be written as
$\mathbf{r}^{\prime}=\mathbf{r}-\mathbf{u}(\mathbf{r})\,,$ (5)
moving points $\mathbf{r}$ (over some domain $\Omega$) in the undeformed ground state to points $\mathbf{r}^{\prime}$ (with some domain $\Omega^{\prime}$). This basis describes the deformed material in the lab frame of an external observer. To clarify, this is considered an active transformation which physically acts on the material and changes the domain of the system in the lab frame. The basis $\mathbf{r}^{\prime}$ is a Cartesian coordinate system in the lab frame of the external observer, but is a non-Cartesian coordinate system in the frame of the material.
We then define an inverse passive transformation which maps the coordinate system $\mathbf{r}^{\prime}$ to the coordinate system $\mathbf{r}^{\prime\prime}$, which is a non-Cartesian basis in the material frame that has the same domain as the inital $\mathbf{r}$ (domain $\Omega$). This passive transformation is defined by the displacement $\mathbf{u}^{\prime}(\mathbf{r}^{\prime})$, where we note that this does not change the energy of the system in any way. It is simply a redefinition of coordinates in order for us to express the actual change in the energy that occured during the active transformation defined in (5).
This passive transformation is given by the form
$\mathbf{r}^{\prime\prime}=\mathbf{r}^{\prime}+\mathbf{u}^{\prime}(\mathbf{r}^{\prime})\,,$ (6)
where the basis $\mathbf{r}^{\prime\prime}$ coincides identically with the basis $\mathbf{r}$ in the material frame, i.e. $\mathbf{r}^{\prime\prime}=\mathbf{r}$, but now the space has additional components in its gradients, as well as a Jacobian associated with its volume element. Note that the passive transformation $\mathbf{u}^{\prime}(\mathbf{r}^{\prime})$ is in the deformed lab frame given by the coordinates $\mathbf{r}^{\prime}$, and importantly is a different transformation to $\mathbf{u}(\mathbf{r})$.
These extra curvatures are necessary to describe the full effect of the displacement procedure on the system. We now relabel the basis $\mathbf{r}^{\prime\prime}$ as $\mathbf{r}$, for convenience, where again we must stress that even though the coordinate systems $\mathbf{r}$ and $\mathbf{r}^{\prime\prime}$ coincide, the space is no longer Cartesian and has an associated Jacobian and curvature. This procedure is illustrated schematically in Figure 2.
Figure 2: A schematic of the deformation procedure. The leftmost image depicts an undeformed system with a basis $\mathbf{r}$, which then undergoes an active transformation given by some strain to become the middle system. The rightmost image depicts the system after the passive transformation, such that the final coordinate system $\mathbf{r}^{\prime\prime}$ coincides exactly with $\mathbf{r}$, but now there are additional curvatures in the space.
As of now, the above transformation is completely general, but in order to proceed we need to specify a particular class of transformation which allows us to use linear response theory and elasticity theory. We make the distinction that $\mathbf{u}({\mathbf{r}})$ is such that all gradients in space are small constants and that gradients in time are considered small velocities.
More specifically, we choose the tensor $\partial_{i}u_{k}$, which is a dimensionless number, s.t. $\partial_{i}u_{k}\sim\delta$, where $\delta\ll 1$. The time derivative of $\mathbf{u}$, i.e. $\partial_{t}u_{k}$, is treated as small in the same sense as the phase perturbation, which we will later cover in Section V. This is referred to as a uniform deformation, in that the strain tensor is constant and small everywhere, which allows us to treat it as a perturbative parameter which we will expand up to harmonic terms.
We now refer to $u_{ik}\equiv\partial_{i}u_{k}$ as the strain tensor, and endeavour to expand $\mathcal{L}$ in powers of $u_{ik}$. In order to employ linear elastic theory we have to keep all terms up to $\mathop{}\!\mathcal{O}(\delta^{2})$ in the Lagrangian. We also need to relate the gradient of the transformation $\mathbf{u}^{\prime}$ in the primed deformed basis in the lab frame to the gradient of the transformation $\mathbf{u}$ in the material basis. It is relatively easy to show that
$\partial_{i}^{\prime}u_{k}^{\prime}=\partial_{i}u_{k}+\partial_{i}u_{l}\partial_{l}u_{k}\equiv u_{ik}+u_{il}u_{lk}$ (7)
up to $\mathop{}\!\mathcal{O}(\delta^{2})$ in the strain tensor (n.b. we are using the Einstein summation convention). This allows us now to express derivatives in the primed (deformed lab frame) basis in terms of derivatives in the unprimed (deformed material frame) basis to find
$\partial_{i}^{\prime}=\partial_{i}+\partial_{k}\left(u_{ik}+u_{il}u_{lk}\right)\quad.$ (8)
In order to simplify the expressions that we will derive in this paper, we introduce notation for a set of key tensors that appear regularly.
$\begin{split}&\epsilon_{ik}=-\frac{1}{2}\left(u_{ik}+u_{ki}\right)\quad;\quad\Delta_{ik}=\frac{1}{2}u_{il}u_{kl}\quad;\\ &\omega_{ik}=\frac{1}{2}u_{li}u_{lk}\quad;\quad\chi_{ik}=\frac{1}{2}\left(u_{li}u_{kl}+u_{lk}u_{il}\right)\end{split}$ (9)
Since the deformation actively changes the domain of the system, there is an associated change of volume. We refer to the deformations described in (5) and (6) to define the differential volume elements as
$\int_{\Omega^{\prime}}\mathop{}\!\mathrm{d}\mathbf{r}^{\prime}=\int_{\Omega}\mathcal{J}_{\mathbf{r}^{\prime}\rightarrow\mathbf{r}}\mathop{}\!\mathrm{d}\mathbf{r}$ (10)
with the Jacobian
$\mathcal{J}_{\mathbf{r}^{\prime}\rightarrow\mathbf{r}}=\det(\frac{\partial r_{i}^{\prime}}{\partial r_{k}})=1+\epsilon_{ll}+M_{iklm}\,u_{ik}u_{lm}\quad.$ (11)
where $M_{iklm}\equiv(\delta_{ik}\delta_{lm}-\delta_{im}\delta_{kl})/2$. A negative $\epsilon_{ll}$ corresponds to a reduction in volume. In two dimensions, $M_{iklm}u_{ik}u_{lm}=u_{xx}u_{yy}-u_{xy}u_{yx}$. The quantities $\epsilon_{ll}$ and $M_{iklm}u_{ik}u_{lm}$ are sometimes called the first and second strain invariants as they are independent of the basis in which the strain tensor is written. We stress now that the coordinate system $\mathbf{r}$ describes the deformed system in the frame of the material: it has the same domain as the undeformed material but is now a non-Cartesian space.
To understand how this deformation affects the Lagrangian, we need to understand how the density changes as a function of the deformation. After the deformation occurs, the particles in the system will reorganise themselves in such a way as to minimise the total energy of the new configuration, leading to a new density described by $\rho(\mathbf{r}^{\prime})$. The reorganisation of particles due to a deformation can be expressed as the pre-deformation density plus a component, $\tilde{\rho}$, which completely accounts for all changes. The component $\tilde{\rho}$ covers both local density changes at length scales within a unit cell of the supersolid and also includes the change in average density due to total number conservation and a change in volume.
We consider that this new density is written in the $\mathbf{r}^{\prime}$ basis. We stress that this is not the density in the material frame, but rather the density in the lab frame of the external observer. We can then use the passive transformation to write the density in the $\mathbf{r}^{\prime\prime}$ basis, which is the material frame that now has some curvature. Without loss of generality we can express this new density at some point $\mathbf{r}^{\prime}$ as
$\begin{split}\rho(\mathbf{r}^{\prime})&=\rho_{0}(\mathbf{r}^{\prime\prime})+\tilde{\rho}(\mathbf{r}^{\prime\prime})\,,\int_{\Omega^{\prime}}\!\!\!\rho(\mathbf{r}^{\prime})\mathop{}\!\mathrm{d}\mathbf{r}^{\prime}=\int_{\Omega}\!\!\!\rho_{0}(\mathbf{r}^{\prime\prime})\mathop{}\!\mathrm{d}\mathbf{r}^{\prime\prime}\end{split}$ (12)
where $\rho_{0}$ is the pre-deformation ground-state density profile. As of now this is simply a mapping to a scalar function that has been subject to an active transformation.
We stress that from this point we drop the $\mathbf{r}^{\prime\prime}$ notation and refer to the coordinate as $\mathbf{r}$ where now the coordinate system is non-Cartesian and has a Jacobian associated with it.
The general form of $\tilde{\rho}$ is determined by minimising the energy of the deformed system, subject to particle conservation. For linear elastic theory, we only need to consider only a perturbative strain on the system and expect that, since the strain is perturbative, the density change $\tilde{\rho}$ must be analytic in the strain tensor. We truncate the Taylor expansion to $\mathop{}\!\mathcal{O}(\delta^{2})$ in the strain tensor and use the ansatz
$\tilde{\rho}(\mathbf{r})=\rho_{1}^{ik}(\mathbf{r})\,\epsilon_{ik}+\rho_{2}^{iklm}(\mathbf{r})\,u_{ik}u_{lm}$ (13)
which we will justify later. This allows us to expand the new normalisation condition for particle conservation.
$\int_{\Omega}(\rho_{0}+\tilde{\rho})\left(1+\epsilon_{ll}+M_{iklm}\,u_{ik}u_{lm}\right)\mathop{}\!\mathrm{d}\mathbf{r}=\int_{\Omega}\rho_{0}\mathop{}\!\mathrm{d}\mathbf{r}$ (14)
and after collecting powers of the strain tensor obtain normalisation requirements for the different terms in the expansion
$\begin{split}\int_{\Omega}\left(\rho_{1}^{ik}+\rho_{0}\delta^{ik}\right)\mathop{}\!\mathrm{d}\mathbf{r}&=0\\ \int_{\Omega}\left[\rho_{2}^{iklm}+\rho_{0}\left(M_{iklm}-\delta_{ik}\delta_{lm}\right)\right]\mathop{}\!\mathrm{d}\mathbf{r}&=0\,.\end{split}$ (15)
Due to the changing energy of the system (and indeed the change in average density), we can expect to see a change in the chemical potential $\mu$ too. In a similar spirit as that of equation (13), we can analytically expand the new chemical potential in the strain tensor as
$\mu=\mu_{0}+\mu_{1}^{ik}\,\epsilon_{ik}+\mu_{2}^{iklm}\,u_{ik}u_{lm}\quad.$ (16)
where the first term is the chemical potential of the undeformed system, and the constants $\mu_{1}^{ik}$ and $\mu_{2}^{iklm}$ encodes how the chemical potential changes due to the deformation and are to be determined by the least-action principle.
We additionally need to consider how the displacement field $u_{ik}$ changes the interaction potential $U$. We note that the interaction is a function of separation between particles, but importantly is originally written in terms of the separation $\Delta r^{\prime}$ as measured in the real lab frame coordinates: $U=U(\Delta r^{\prime})$ For a uniform strain, we find immediately that the lab-frame separation $\Delta\mathbf{r}^{\prime}$ can be written in terms of the material-frame separation $\Delta\mathbf{r}$ such that
$\left(\Delta r^{\prime}\right)^{2}=\left(\delta_{ik}+2\epsilon_{ik}+2\Delta_{ik}\right)\left(\Delta r\right)_{i}\left(\Delta r\right)_{k}$ (17)
which is exactly equivalent to that of Landau & Lifshitz [28]. The tensor in the above equation is sometimes referred to as the finite strain tensor. A Taylor expansion of the interaction for small strain gives
$\begin{split}U(\Delta\mathbf{r}^{\prime})&=U(\Delta\mathbf{r})+\left(\epsilon_{ik}+\Delta_{ik}\right)f_{ik}(\Delta\mathbf{r})\\ &\quad+\epsilon_{ik}\epsilon_{lm}W_{iklm}(\Delta\mathbf{r})\end{split}$ (18)
with
$\begin{split}f_{ik}(\mathbf{r})&\equiv\frac{r_{i}r_{k}}{|\mathbf{r}|}\frac{\partial U(|\mathbf{r}|)}{\partial|\mathbf{r}|}\\ W_{iklm}(\mathbf{r})&\equiv\frac{r_{i}r_{k}r_{l}r_{m}}{2|\mathbf{r}|^{2}}\left(\frac{\partial^{2}U}{\partial|\mathbf{r^{2}}|}-\frac{1}{|\mathbf{r}|}\frac{\partial U}{\partial|\mathbf{r}|}\right)\;.\end{split}$ (19)
To summarise, we have now defined a perturbative deformation which changes the geometry of our system, and have obtained the volume changes and curvature changes due to said deformation. We have also derived a new density and chemical potential, both of which we will solve for using the least-action principle. We now have all the ingredients to formulate a theory of linear response of the system to elastic deformations.
IV Elasticity Theory
We can now develop a linear elastic theory by expanding the Lagrangian to second order in the strain tensor and finding the response, $\tilde{\rho}$, of the density to a small applied strain using the principle of least action. The full details are given in Appendix A where we calculated and collected all the first- and second-order components of $u_{ik}$ in the Lagrangian. This should be of the form:
$L(\Omega^{\prime})\simeq L_{0}(\Omega)+|\Omega|\pi_{ik}u_{ik}-\frac{|\Omega|}{2}A_{iklm}u_{ik}u_{lm}$ (20)
where the first term is the ground-state Lagrangian, the second term contains the coupling to the stress tensor $\pi_{ik}$, and the third term is related to the Cauchy elastic tensor $B_{iklm}$, as carefully discussed in Bavaud et al. [29]. ($|\Omega|$ is the total volume of the pre-deformed system.)
Let us consider the term linear in the strain tensor. The pressure tensor is related to the stress tensor $\pi_{ik}$ by $P_{ik}=-\pi_{ik}$. We will be studying a two-dimensional supersolid whose spatial modulation forms a triangular lattice. The $C_{3}$ symmetry of the ground state dictates that $P_{ik}=P\delta_{ik}$, i.e. the pressure is isotropic. Collecting the $\mathop{}\!\mathcal{O}(u_{ik})$ terms in the expansion of the Lagrangian [see (75) and (76)], we can identify the pressure as
$\begin{split}P_{ik}=\int_{\Omega}&\left[\frac{\hbar^{2}}{4m}\left(\frac{\partial_{i}\rho_{0}\partial_{k}\rho_{0}}{\rho_{0}}-\nabla^{2}\rho_{0}\,\delta_{ik}\right)\right.\\ &\left.\qquad\qquad\qquad-\frac{1}{2}\rho_{0}(f_{ik}*\rho_{0})\right]\frac{\mathop{}\!\mathrm{d}\mathbf{r}}{|\Omega|}\end{split}$ (21)
which will be used in subsequent calculations. The above expression can be directly calculated for any ground state density $\rho_{0}$. In the case of a spatially homogeneous superfluid, it leads to what we expect: $P_{ik}=\frac{1}{2}\mu_{0}n\delta_{ik}$. (See Appendix A.)
Let us now turn to the the second-order terms in the expansion (20). It is important to note that the tensor $A_{iklm}$ is not the Cauchy elastic tensor $B_{iklm}$ which provides the tensor in the Hooke’s law for the stress/strain relationship between a deformed state stress and the strain that induced the deformation, i.e. $\pi_{ik}(\Omega^{\prime})=\pi_{ik}(\Omega)-B_{iklm}(\Omega)u_{lm}$. It can be shown [29] that:
$B_{iklm}=A_{iklm}+P_{ik}\delta_{lm}-P_{im}\delta_{kl}$ (22)
which we will derive independently later in (VI) by considering work done by the expansion on the surroundings.
For a supersolid with $C_{3}$ symmetry, we also expect the elastic tensor to obey the indicial symmetries $B_{iklm}=B_{klim}=B_{lkmi}=B_{mlki}$ so that it only contains two independent quantities — the bulk modulus $K=c_{44}$ and the shear modulus $G=c_{66}$:
$B_{iklm}=K\delta_{ik}\delta_{lm}+G\left(\delta_{il}\delta_{km}+\delta_{im}\delta_{kl}\right)\,.$ (23)
It should be noted that $A_{iklm}$ is not indicially symmetric in the same way as $B_{iklm}$. Elastic theory [28] in 2D posits that the bulk modulus $K$ is given by $B_{xxyy}$, and the shear modulus $G$ is given by $B_{xyxy}$ (or all equivalent indicially symmetric elements). For a triangular lattice or a homogeneous system, the tensor also possesses the symmetry: $B_{xxxx}=B_{xxyy}+2B_{xyxy}$. Note also that we expect the changes to the chemical potential to have the same symmetry as the pressure, i.e. $\mu_{1}^{ik}=\mu_{1}\delta_{ik}$. We will later make use of these properties to simplify the expressions for the elastic constants.
We can deduce the elastic tensor by identifying the $\mathop{}\!\mathcal{O}(u_{ik}^{2})$ terms in the expansion of the Lagrangian with $A_{iklm}$, and then using (22) to obtain $B_{iklm}$. (See Appendix (88) and the derivative rules in Appendix (C)). We find that $B_{iklm}=\absolutevalue{\Omega}^{-1}\int_{\Omega}b_{iklm}(\mathbf{r})\mathop{}\!\mathrm{d}\mathbf{r}$ where
$\begin{split}b_{iklm}(\mathbf{r})=&+\frac{\hbar^{2}}{4m}\left(\frac{\partial_{i}\rho_{0}\partial_{l}\rho_{0}}{\rho_{0}}\delta_{km}+\frac{\partial_{k}\rho_{0}\partial_{l}\rho_{0}}{\rho_{0}}\delta_{im}-\frac{\partial_{i}\rho_{0}\partial_{k}\rho_{0}}{\rho_{0}}\delta_{lm}\right)\\ &+\frac{1}{2}\left(f_{il}*\rho_{0}\right)\rho_{0}\delta_{km}-\frac{1}{2}\left(f_{ik}*\rho_{0}\right)\rho_{0}\delta_{lm}+\frac{1}{2}\left(f_{im}*\rho_{0}\right)\rho_{0}\delta_{kl}\\ &+\left[\left(\left(U\delta_{ik}+2f_{ik}\right)\delta_{lm}+W_{iklm}\right)*\rho_{0}\right]\rho_{0}+\rho_{1}^{lm}\left[\frac{\hbar^{2}}{4m}\left(2\frac{\partial_{ik}\rho_{0}}{\rho_{0}}-\frac{\partial_{i}\rho_{0}\partial_{k}\rho_{0}}{\rho_{0}^{2}}\right)+\left(f_{ik}+\delta_{ik}U\right)*\rho_{0}+\mu_{1}^{ik}\right]\end{split}$ (24)
with $(g*\rho)_{\mathbf{r}}\equiv\int\mathop{}\!\mathrm{d}\mathbf{r}^{\prime}g(\mathbf{r}-\mathbf{r}^{\prime})\rho(\mathbf{r}^{\prime})$ for any function $g$ and $\rho$ and where $f_{ik}$ and $W_{iklm}$ are defined in (19).
We note that this result for the elastic constants depends on the ground state density $\rho_{0}$ as well as the shift in the density $\rho_{1}$ and shift in the chemical potential $\mu_{1}$ as defined in (13) and (16). In other words, we need to deduce how the density profile has changed in response to the applied strain. This is achieved by solving the Euler-Lagrange equations for the variable $\tilde{\rho}$ as defined in (12). This involves solving the following equation
$\begin{split}&\frac{\hbar^{2}}{4m}\gradient\cdot\left(\frac{\gradient\tilde{\rho}}{\rho_{0}}\right)-U*\tilde{\rho}\\[5.0pt] +&\frac{\hbar^{2}}{4m}\left(\frac{(\nabla\rho_{0})^{2}}{\rho_{0}^{3}}-\frac{\nabla^{2}\rho_{0}}{\rho_{0}^{2}}\right)\tilde{\rho}\end{split}\text{\Large{\hbox{ = }}}\begin{split}&-\left(\mu_{2}^{iklm}+\mu_{1}^{ik}\delta_{lm}+\mu_{0}M_{iklm}\right)u_{ik}u_{lm}+\epsilon_{ik}\left[-\mu_{1}^{ik}+\frac{\hbar^{2}}{4m}\left(2\frac{\partial_{ik}\rho_{0}}{\rho_{0}}-\frac{\partial_{i}\rho_{0}\partial_{k}\rho_{0}}{\rho_{0}^{2}}\right)\right.\\[5.0pt] &\left.+\int_{\Omega}\left(f_{ik}(\mathbf{r}-\mathbf{r}^{\prime})+\delta_{ik}U(\mathbf{r}-\mathbf{r}^{\prime})\right)\rho_{0}(\mathbf{r}^{\prime})\mathop{}\!\mathrm{d}\mathbf{r}^{\prime}\right]\end{split}$ (25)
which is an integro-differential equation that can be solved either by using Bloch’s theorem (only applicable in periodic boundary conditions) or solving directly via writing the linear operator as a matrix and using standard iterative matrix solvers. More details on solving the equation and finding the chemical potentials are provided in Appendix (79).
It must be noted that we actually have 8 equations to solve, as for each pair of indices $(i,k)$ there is a different RHS source function, and for each of these we need to solve the appropriate equation for both $\rho_{1}^{ik}$ and $\mu_{1}^{ik}$. This generates 8 equations for a 2D system and 18 equations for a 3D system. Once we have the solutions for both $\rho_{1}^{ik}$ and $\mu_{1}^{ik}$, we can calculate the other functions in the expression for (24) and obtain the elastic tensor $B_{iklm}$.
Let us check these results for the case of a homogeneous system, where the modulation vanishes and we can check our results against the known results for elastic moduli of a superfluid. We can see that the elastic tensor is greatly simplified by the disappearance of the derivative terms $\partial_{i}\rho_{0}$, and the chemical potential is analytically known: $\mu=U*\rho_{0}$. We can find the convolutions $f_{ik}*\rho_{0}$ and $W_{iklm}*\rho_{0}$, solve the equation (79) and use this to find that
$B_{iklm}=\mu n\delta_{ik}\delta_{lm}\quad.$ (26)
Using the definitions of $K$ and $G$ from the Cauchy elastic tensors (23), we see here that we obtain the expected results of $K=\mu n$ and no shear modulus: $G=0$.
We will now present the numerical calculation of these homogenisation results for the elastic constants. The results are shown in Figure 3 (solid markers). In our numerics, we see some small asymmetry, e.g. $B_{iklm}\approxeq B_{lmik}$, in the results for the elastic tensor. The discrepancy with the ideal form (23) decreases as we reduce the discretisation and increase the system size in the numerics. With that in mind, we choose to define the bulk modulus $K$ as the average of the values of $B_{xxyy}$ and $B_{yyxx}$, the pressure $P$ as the average of $P_{xx}$ and $P_{yy}$, and the shear modulus $G$ as the average of $B_{xyxy}$, $B_{xyyx}$, $B_{yxyx}$ and $B_{yxxy}$. We then use the minimal and maximal values that the elastic moduli could have taken in order to generate ‘error bars’ for the elastic constants. It should be stressed that these error bars are not measures of some statistical error, but rather a quantification of how much the elastic constants deviate due to the crystal symmetry being only approximately triangular (as opposed to exactly triangular) as a result of discretisation error in our numerics.
Figure 3: A comparison of the various elastic constants obtained via homogenisation to those from numerics (homogenisation constants are shown in solid markers, numerics in crosses). The top figure shows the bulk modulus and shear modulus, the middle figure shows the pressure, and the bottom figure shows the absolute value of $\mu_{1}$. The agreement between homogenisation and numerics is very strong on the whole, but falls apart in the phase transition region. This is likely due to numerical instabilities close to the phase transition which affect the numerically obtained constants but not the homogenisation constants.
We now turn to comparing these numerical results from homogenisation with a direct calculation of the elastic constants by finding how the energy of the ground state of the Gross-Pitaevskii equation changes when we deform the simulation cell at fixed particle number (with periodic boundary conditions). We begin with an undeformed ground-state obtained via numerically solving the Gross-Pitaevskii with some given average density, area and interparticle interaction. We then apply a small strain to the system by changing the geometry of the simulation cell by a strain tensor $u_{ik}$, whilst keeping the total number of particles fixed, and solving the Gross-Pitaevskii equation for the new geometry to find the deformed ground-state. For sufficiently small strain $u_{ik}$, we expect that such a procedure will yield a ground state with a Lagrangian of the form (20), where the tensor $u_{ik}$ is now a control parameter in our computational experiment. This strain is small in the same sense as defined in the expansion of the Lagrangian. More precisely, we consider $u_{ik}\ll 1$, and then try to find the harmonic terms in the energy perturbation.
By numerically extracting the Lagrangian for a system which has been deformed by some strain $u_{ik}$, we are able to find the pressure and elastic constants by judicious selection of strains and subsequent fitting of the coefficients. The full procedure is described in Appendix D, where we show the extraction of elastic constants specifically for a 2D case in a triangular geometry.
Our results are presented in Figure 3, where we have plotted the bulk and shear modulus $K$ and $G$, the pressure $P$ and the shift in the chemical potential for unit strain: $\mu_{1}^{ik}=\mu_{1}\delta_{ik}$. These elastic constants should ideally be the same as those obtained via homogenisation theory. In reality, they suffer from numerical errors due to being unable to perform an arbitrarily small strain, as well as convergence errors in the new ground-state when close to the first-order phase transition between the homogeneous superfluid and the supersolid. Scanning across the whole range of the interaction parameter $\Lambda$, we find that, outside of a small region where the phase transition occurs, the elastic constants obtained numerically are in excellent agreement with those obtained analytically, with a typical discrepancy of $\sim 0.1\%$.
We found that an applied numerical strain which was smaller than $\sim 5\%$ (by this we mean $u_{ik}\sim 0.05$) would generate this excellent agreement, but if the applied strain was $\sim 10\%$, the numerical value would diverge from that of homogenisation. We suspect that this is due to the introduction of anharmonic elastic moduli which are now no longer negligible.
In summary, we have developed an elastic theory for a supersolid using homogeneration theory and calculated its elastic properties using the solution of the Gross-Pitaevskii equation. Our homogenisation results agree well with results from the numerical simulation of a strained supersolid.
V $U(1)$ Phase and Fractional Inertia
In this section, we consider the application of a phase perturbation to the meanfield supersolid, and analyse the coupling of the phase to the strain field. We follow the metholodology of Josserand et al [15], but our results have some key differences that we will point out.
We begin by specifying the nature of the phase perturbation. We consider imposing a small superflow to the system in the form of a phase gradient. This is applied to the supersolid in the deformed state, i.e the lab frame of the deformed material. The system will respond to this long-wavelength perturbation by generating phase fluctuations, $\tilde{\phi}$, at small wavelengths within a unit cell of the supersolid. As with $\tilde{\rho}$ in the previous section, $\tilde{\phi}$ is a correcting function to the respective perturbation applied to the ground state. Namely, $\tilde{\rho}$ is the function which allows the density to settle into the energetically optimal configuration for a given strain, and $\tilde{\phi}$ is the function which will allow the phase field to settle into the energetically optimal configuration for a given applied phase gradient. We mathematically represent this as
$\phi(\mathbf{r}^{\prime})=\phi_{0}(\mathbf{r}^{\prime})+\tilde{\phi}(\mathbf{r})$ (27)
where $\phi(\mathbf{r}^{\prime})$ is the total phase in the lab frame $\mathbf{r}^{\prime}$, $\phi_{0}(\mathbf{r}^{\prime})$ is the applied perturbative phase in the lab frame, and $\tilde{\phi}(\mathbf{r})$ is the phase-correction provided by the material in the material frame of reference. We express $\tilde{\phi}$ in terms of the material frame deformed coordinate system $\mathbf{r}$ in keeping with linear response theory, similar to how we expressed $\tilde{\rho}$ in terms of the material frame coordinate system $\mathbf{r}$.
In order to apply perturbation theory to the phase field, we specify that $\partial_{i}\phi_{0}(\mathbf{r}^{\prime})$ is small and constant within the scope of the unit cell of the supersolid, in a similar way to the strain tensor $u_{ik}$. This is equivalent to stating that the phase change across the unit cell is much smaller than $2\pi$, i.e. $\Delta\phi\ll 2\pi$. Note also that we are not considering the generation of vortices.
We note that since the phase is applied in the deformed frame, the total phase in the material frame can be expanded to obtain
$\phi(\mathbf{r}^{\prime})=\phi_{0}(\mathbf{r})-\mathbf{u}\cdot\gradient\phi_{0}(\mathbf{r})+\tilde{\phi}(\mathbf{r})$ (28)
where $\mathbf{u}$ is the displacement as defined in (5). This is the final expression for the phase field in the material frame of reference. We can now proceed to expand the Lagrangian in terms of the phase field, using the least action principle to find $\tilde{\phi}$ and the associated energy.
Using the transformation rules for the Lagrangian (Jacobian, derivatives etc.) and the $\phi$ expansion given by (28), the phase-dependent part of the Lagrangian (1) can be expanded to $\mathop{}\!\mathcal{O}(\tilde{\phi}^{2})$ as
$\begin{split}\mathcal{L}_{\phi}=&-\int_{\Omega}\Big{\{}\hbar\rho_{0}\partial_{t}\phi_{0}\left(1+\gradient\cdot\mathbf{u}\right)\\ &+\frac{\hbar^{2}\rho_{0}}{2m}\left[\left(\gradient\phi_{0}\right)^{2}+2\mathbf{A}\cdot\gradient\tilde{\phi}+(\gradient\tilde{\phi})^{2}\right]\Big{\}}\mathop{}\!\mathrm{d}\mathbf{r}\end{split}$ (29)
where
$\mathbf{A}=\gradient\phi_{0}+\left(\gradient\phi_{0}\cdot\gradient\right)\mathbf{u}+\frac{m}{\hbar}\partial_{t}\mathbf{u}\,.$ (30)
The vector field $\mathbf{A}$ is a kind of convective flow. First of all, recall that the superfluid velocity is given by $\mathbf{v}=(\hbar/m)\gradient{\phi_{0}}$. Also, we can define the material derivative of a field as
$\frac{D}{Dt}=\partial_{t}+\mathbf{v}\cdot\gradient$ (31)
to describe the time derivative of a function which is co-moving with the superfluid velocity field.
Note here the appearance of the term $m/\hbar\partial_{t}\mathbf{u}$ which we previously stated as ‘small’ without further elaboration. Here we define this term to be small on the same order as $\mathbf{\nabla}\phi_{0}$, i.e. that the velocity induced by the elastic deformation is comparable to the velocity induced by the phase change.
We can rewrite $\mathbf{A}$ in term of these variables:
$\mathbf{A}=\frac{m}{\hbar}\left(\mathbf{v}+\mathbf{v}\cdot\gradient\mathbf{u}+\partial_{t}\mathbf{u}\right)=\frac{m}{\hbar}\left(\mathbf{v}+\frac{D\mathbf{u}}{Dt}\right)\,.$ (32)
A point of subtlety is that the actual deformation we have applied is $\mathbf{r}^{\prime}=\mathbf{r}-\mathbf{u}$, and so $(\hbar/m)\mathbf{A}=\mathbf{v}-D(-\mathbf{u})/Dt$ is a relative velocity between the superflow and the motion of the lattice deformations. This is exactly the form predicted by Son [12] from symmetry based arguments. In fact, all terms in $\mathcal{L}_{\phi}$ satisfy the Galilean symmetry requirements imposed by Son. In this way we can show that we have microscopically satisfied the necessary Galilean symmetry requirements imposed on any supersolid system macroscopically.
Our phase Lagrangian $\mathcal{L}_{\phi}$ in (29) differs from the analogous form in Josserand et al [15] by the existence of the term $\hbar\rho_{0}\partial_{t}\phi_{0}\gradient\cdot\mathbf{u}$. This term is in fact necessary to ensure that $\mathcal{L}_{\phi}$ satisfies Galilean symmetry constraints as pointed out by Son [12]. We also note that this affects the velocity of the excitations of the system, specifically in reference to (46) where this term is directly responsible for the $n$ component o $\hbar\left(n-\varrho+\mu_{1}/\mathcal{E}^{\prime\prime}\right)\dot{\phi}\mathbf{\nabla}\cdot\mathbf{u}$. If one we re to calculate th velocities without this critical term, one would see that it differs significantly from those predicted by Bogoliubov theory (shown in Section VIII).
Figure 4: The superfluid fraction $f_{s}\equiv 1-\varrho/n$ as a function of $\Lambda$. We appear to observe a discontinuous transition at $\Lambda\sim 38$, which remains even after checking the transition region with a much finer $\Lambda$ discretisation. This appears to be numerical verification of [17], which predicted a first order transition close to this value.
Let us now solve for the short-distance phase response $\tilde{\phi}$ within a unit cell due to superflow and strain. We follow a similar procedure to Ref [15]. By recognising that the externally imposed phase twist and strains $\gradient\phi_{0}$ and $\partial_{i}u_{k}$ are defined as long-wavelength relative to the unit cell, we can treat $\mathbf{A}$ as constant within a unit cell. However, a uniform velocity with non-uniform density does not necessarily obey local mass conservation in equilibrium. So we need a correction to the phase field, $\tilde{\phi}$, to ensure that our system obeys the continuity equation in the ground state: $\gradient\cdot\left(\rho\gradient\phi\right)=0$. The Euler-Lagrange equation for $\tilde{\phi}$ gives
$\tilde{\phi}=K_{i}A_{i}\quad\mbox{with}\quad\partial_{i}\rho+\mathbf{\gradient}\cdot\left(\rho\mathbf{\gradient}K_{i}\right)=0\,.$ (33)
This can be solved numerically in much the same way as (80). Physically, this result expresses that the fact the phase twist (and therefore superflow) is not uniform within a unit cell because it costs less kinetic energy to introduce phase shifts in regions of low density compared to regions of high density. Using this solution for $\tilde{\phi}$, it can be shown that the phase-dependent Lagrangian can be written as:
$\begin{split}L_{\phi}=&\absolutevalue{\Omega}\frac{\hbar^{2}}{2m}\varrho_{ij}A_{i}A_{j}-\int_{\Omega}\Big{[}\hbar\rho_{0}\partial_{t}\phi_{0}\left(1+\gradient\cdot\mathbf{u}\right)\\ &\qquad\qquad\qquad\qquad+\frac{\hbar^{2}\rho_{0}}{2m}\left(\gradient\phi_{0}\right)^{2}\Big{]}\mathop{}\!\mathrm{d}\mathbf{r}\end{split}$ (34)
with
$\varrho_{ij}=-\frac{1}{\absolutevalue{\Omega}}\int_{\Omega}\rho_{0}\partial_{i}K_{j}\mathop{}\!\mathrm{d}\mathbf{r}$ (35)
which can be interpreted as a correction to the superfluid phase stiffness (see equation (37)). Figure 4 shows the evolution of this quantity as we increase the interaction strength $\Lambda$ showing a discontinuous transition at the superfluid-supersolid transition.
To restate the above, by carefully considering the dynamics of both an applied/induced phase and a strain deformation (both of which are long-wavelength relative to the unit cell), we calculate how the phase-dependent dynamics of our system are dependent on the parameters $n$ and $\varrho_{ij}$ given by the form (34). $n$ is simply the density of particles in the system, and corresponds to the phase-coherent BEC with components at both $\mathbf{k}=0$ and $\mathbf{k}\neq 0$. We will later show that $\varrho_{ij}$ is the part of the system that corresponds to the pattern formation, i.e. it only exists when $\psi_{\mathbf{k}\neq 0}\neq 0$. Consequently, we refer to $\varrho_{ij}$ as being the “supersolid fraction” of the system. In this way we have separated out the pattern response of the system to the coupled phase-strain dynamical field from the standard superfluid response.
Finally, due to the $C_{3}$ symmetry, second-rank tensors representing bulk properties should be isotropic. We therefore make the assumption that $\varrho_{ij}=\varrho\delta_{ij}$, though in principle this is not necessary. We also note that the dependency in (34) on $\phi_{0}$ and $\mathbf{u}$ can be taken out of the integral (due to the same argument that $\mathbf{A}$ is a constant in the unit cell). Combining these two simplifications, we find that we can write
$\begin{split}\frac{L_{\phi}}{|\Omega|}=&-\hbar n\dot{\phi}_{0}(1+\gradient\cdot\mathbf{u})\\ &-\frac{\hbar^{2}}{2m}\left[n\left(\gradient\phi_{0}\right)^{2}-\varrho\left(\gradient\phi_{0}+\frac{m}{\hbar}\frac{D\mathbf{u}}{Dt}\right)^{2}\right]\end{split}$ (36)
where $n=\int_{\Omega}\rho_{0}\mathop{}\!\mathrm{d}\mathbf{r}/|\Omega|$ is the average particle density. We use this result, along with the elastic expansion (24), to write out the full Lagrangian after the homogenisation process as
$\frac{L_{\mathrm{hom}}}{\absolutevalue{\Omega}}=L_{0}-\hbar n\dot{\phi}_{0}(1+\gradient\cdot\mathbf{u})-P_{ik}u_{ik}-\frac{1}{2}A_{iklm}u_{ik}u_{lm}-\frac{\hbar^{2}}{2m}\left(n-\varrho\right)\left(\gradient\phi_{0}\right)^{2}+\hbar\varrho\gradient\phi_{0}\cdot\frac{D\mathbf{u}}{Dt}+\frac{1}{2}m\varrho\left(\frac{D\mathbf{u}}{Dt}\right)^{2}$ (37)
This effective Lagrangian contains all the linear response of the system due to both strain and phase gradients. We can see that the phase stiffness is now given by $n-\varrho$ and $\varrho$ is the inertia associated with the motion of the modulated pattern of the density. In summary, we have followed Ref. [15] and reproduced the same results for the superfluid phase stiffness of the supersolid. Our Lagrangian differs from that reference in the dynamical term. In the next section, we complete the development of an effective low-energy theory for the supersolid by considering the free energy of the system in the grand canonical ensemble. Then, we can finally proceed to discuss the macroscopic dynamics of the system.
VI Free energy and Grand Canonical ensemble
In Sections III to V, we developed a homogenised theory of the supersolid in the canonical ensemble at a fixed average particle density of $n$. We have integrated out the intra-unit-cell dynamics of the supersolid and have now obtained an effective Lagrangian $L_{\mathrm{hom}}(\phi_{0},\mathbf{u})$ which describes the system under a given phase gradient and displacements $\mathbf{u}$. We have assumed that all the unit cells responded in the same way, i.e. the response has the same periodicity as the supersolid. This effective Lagrangian obeys all the necessary requirements of Galilean invariance, rotational invariance and $U(1)$ invariance as specified by Son [12] and Andreev & Lifshitz [7], so we would expect the dynamics to be phenomenologically correct.
In this section, we will consider the inter-cell dynamics and allow for particle flow between unit cells leading to variations in the density $\delta n$ across the system at length scales much larger than the size of the unit cells of the supersolid.
It is important to note that the additional energy response due to flow between cells is crucial to obtaining the correct dynamics of the system. If we were to neglect this energy response, we would find that the dynamics of the system would not match up with a Bogoliubov treatment. A similar phenomena was observed in [30], where the authors found that the bulk modulus computed via a homogeneous dilation technique was not in agreement with long-wavelength phonon predictions unless charge redistribution was taken into account.
We assume that the inter-cell number fluctuation is small, i.e. that $\delta n/n\ll 1$, and so we can expand the Lagrangian analytically in $\delta n$ up to $\mathop{}\!\mathcal{O}(\delta n^{2})$. We now consider the ground state Lagrangian density $\mathcal{E}(n)=-L_{0}/|\Omega|$ which is a function of the average density $n$ of the undeformed system. Under a change of the local density, this can be expanded as
$\mathcal{E}(n+\delta n)\simeq\mathcal{E}(n)+\mathcal{E}^{\prime}\delta n+\frac{1}{2}\mathcal{E}^{\prime\prime}(\delta n)^{2}$ (38)
and similarly the local pressure can be expanded as
$P(n+\delta n)\simeq P(n)+P^{\prime}\delta n\,.$ (39)
These are the only terms we need to track in a theory that is an expansion up to second order in the total power of all perturbative fields. The coefficients, $P^{\prime}$ and $\mathcal{E}^{\prime\prime}$, can be calculated numerically by solving Gross-Pitaevskii equation at several densities. We can also use $\mathcal{E}^{\prime}=\mu_{0}$ and ′′$\mathcal{E}^{\prime\prime}=\partial\mu/\partial n$ if we have the density dependence of the chemical potential. In fact, we have already calculated the pressure derivative in terms of the chemical potential shift per unit strain, $\mu_{1}$, as defined in (16) with $\mu_{1}^{ik}=\mu_{1}\delta_{ik}$. To see this, we note that the pressure, $P=-(\partial E/\partial V)_{N}$, and the chemical potential, $\mu=(\partial E/\partial N)_{V}$, are related by a Maxwell relation so that
$P^{\prime}=\frac{\partial P}{\partial n}=V\left(\frac{\partial P}{\partial N}\right)_{V}\!\!\!=-V\left(\frac{\partial\mu}{\partial V}\right)_{N}\!\!\!=-\mu_{1}\,.$ (40)
since $\delta V/V=u_{kk}$ is the strain. Similarly, we can use ′′$\mathcal{E}^{\prime\prime}=\partial\mu/\partial n$ if we have already obtained the density dependence of the chemical potential.
In general, the above analysis is valid for any tensorial quantity, but we have elected to use the symmetry properties of tensors that belong to a triangular lattice (i.e. $P_{ik}=P\delta_{ik}$ and $\mu_{1}^{ik}=\mu_{1}\delta_{ik}$) to simplify our calculations. The following analysis also uses this symmetry, but can be generalised to any tensorial quantity.
Finally, we need to include the work done by the change in local density on the local environment. Namely, if a part of the system expands due to an increase in particle number, it does work on the surroundings which must compress. The work done caused by a total volume change of $\delta V$ is:
$\delta W=-P\delta V=-P(\epsilon_{ll}+M_{iklm}u_{ik}u_{lm})|\Omega|\,.$ (41)
In order to describe the dynamics of these density functions, we need to add this work done to the Lagrangian to obtain
[…]
Comparing with our form (37) for $L_{\mathrm{hom}}$ and the relation (22) between the Cauchy tensor $B_{iklm}$ and $A_{iklm}$, we find that
-- $Pu_{ll}-\frac{A_{iklm}}{2}u_{ik}u_{lm}\!-\!P(\epsilon_{ll}+M_{iklm}u_{ik}u_{lm})$
$\simeq-\frac{1}{2}B_{iklm}u_{ik}u_{lm}\,.$ (43)
Using this, we arrive at an effective quadratic Lagrangian that can describe the dynamics of the low-energy excitations of the supersolid. This is the principal result of our paper:
$\begin{split}\frac{L_{\mathrm{dyn}}}{\absolutevalue{\Omega}}=&-\mathcal{E}-\mu_{0}\delta n-\frac{1}{2}\mathcal{E}^{\prime\prime}(\delta n)^{2}-\hbar n\dot{\phi}_{0}\left(1+\gradient\cdot\mathbf{u}\right)-\hbar\delta n\dot{\phi}_{0}+\mu_{1}\delta n\gradient\cdot\mathbf{u}\\ &-\frac{1}{2}B_{iklm}u_{ik}u_{lm}-\frac{\hbar^{2}}{2m}\left(n-\varrho\right)\left(\gradient\phi_{0}\right)^{2}+\hbar\varrho\gradient\phi_{0}\cdot\dot{\mathbf{u}}+\frac{1}{2}m\varrho\dot{\mathbf{u}}^{2}\,.\end{split}$ (44)
In this process we note that we have directly identified the Cauchy elastic tensor in two different ways, firstly by using the result from Bavaud [29], and secondly by directly considering the work the excitations would need to do on their environment.
To summarise, we have a coarse-grained Lagrangian to describe the low-energy excitations of the supersolid. In the next section, we will investigate these dynamics.
VII Effective supersolid Lagrangian
In the previous sections, we have derived an effective Lagrangian (44) by integrating out the short-wavelength modes in the unit cell and coarse-graining the system such that a unit cell is now considered a point in the continuous field theory described by the fields $\{\mathbf{u},n,\phi\}$. This Lagrangian formally describes a low-energy effective field theory wherein the coupled Goldstone modes of the system are the low-lying excitations. We now set out to calculate the normal modes by integrating out the field $\delta N$ and solving the resulting set of coupled equations.
By minimising the Lagrangian for $\delta n$, we find
$\delta n=\frac{\mu_{1}}{\mathcal{E}^{\prime\prime}}\gradient\cdot\mathbf{u}-\frac{\hbar}{\mathcal{E}^{\prime\prime}}\dot{\phi}$ (45)
which we substitute back into the Lagrangian to straightforwardly obtain
$\begin{split}L=&-\mathcal{E}-\hbar n\dot{\phi}+\frac{\hbar^{2}}{2}\left[\frac{\dot{\phi}^{2}}{\mathcal{E}^{\prime\prime}}-\frac{n-\varrho}{m}\left(\gradient\phi\right)^{2}\right]\\ &+\frac{1}{2}\left[m\varrho\dot{\mathbf{u}}^{2}-\left(B_{iklm}-\frac{\mu_{1}}{\mathcal{E}^{\prime\prime}}\mu_{1}\delta_{ik}\delta_{lm}\right)u_{ik}u_{lm}\right]\\ &-\hbar\left(n-\varrho+\frac{\mu_{1}}{\mathcal{E}^{\prime\prime}}\right)\dot{\phi}\gradient\cdot\mathbf{u}\end{split}$ (46)
This Lagrangian should be able to describe the Goldstone modes for the system. This effective Lagrangian is of the form predicted by Ye [31].
By counting the number of spontaneously broken continuous symmetries, we are able to predict the number of Goldstone modes that emerge. Since we have broken the $U(1)$ gauge symmetry, we expect to find one Goldstone mode. We also have a triangular lattice, which means we have broken two continuous translational and rotational symmetries, and so we expect to find two Goldstone modes. Therefore we should find three Goldstone modes in total.
One can immediately find the equations of motion for both $\phi$ and $\mathbf{u}$, but it serves as a useful check to ensure we recover the Lagrangian for a Bogoliubov mode in the superfluid phase. In a superfluid phase, one can show that $B_{iklm}=U_{k=0}n^{2}\delta_{ik}\delta_{lm}$, $\mu_{1}=-U_{k=0}n$, ′′$\mathcal{E}^{\prime\prime}=U_{k=0}$ and $\varrho=0$. Consequently, the resultant Lagrangian is now
$L=-\mathcal{E}-\hbar n\dot{\phi}+\frac{\hbar^{2}}{2}\left[\frac{\dot{\phi}^{2}}{U_{k=0}}-\frac{n}{m}\left(\gradient\phi\right)^{2}\right]$ (47)
which corresponds exactly with the superfluid Bogoliubov mode.
Note that for the following derivation of the supersolid Goldstone velocities, it is assumed that the ground state is a supersolid and hence $\varrho\neq 0$. In a superfluid state ($\varrho=0$), the following analysis does not apply, and instead one needs to refer to (47).
For the general supersolid case, we derive the equations of motion directly by assuming the fields $\phi$ and $\mathbf{u}$ are of a plane wave form $\sim e^{i\left(\mathbf{k}\cdot\mathbf{r}-\omega t\right)}$. The vector equation for the variable $\mathbf{u}$ can be solved to provide both the shear and longitudinal mode. The Euler-Lagrange equation for the variable $\mathbf{u}$ is immediately found to be
$\begin{split}m\varrho\ddot{\mathbf{u}}-&\left(K+G-\frac{\mu_{1}}{\mathcal{E}^{\prime\prime}}\mu_{1}\right)\gradient\left(\gradient\cdot\mathbf{u}\right)-G\nabla^{2}\mathbf{u}\\ &-\hbar\left(n-\varrho+\frac{\mu_{1}}{\mathcal{E}^{\prime\prime}}\right)\gradient\dot{\phi}=0\end{split}$ (48)
where we have used (23) as we are analysing a triangular lattice.
We can now solve this equation by noting that the vector field $\mathbf{u}$ can be split into longitudinal and transverse elements $\mathbf{u}=\mathbf{u}_{l}+\mathbf{u}_{t}$. Taking the curl of (48) isolates the transverse motion ($\mathbf{u}_{t}$) giving:
×$\gradient\times\left(m\varrho\ddot{\mathbf{u}}_{t}-G\nabla^{2}\mathbf{u}_{t}\right)=0$ (49)
This is a wave equation for ×$\gradient\times\mathbf{u}$ from which we obtain the shear velocity
$c_{\textrm{shear}}=\sqrt{\frac{G}{m\varrho}}$ (50)
which is exactly equivalent to the shear velocity predicted in standard solid elasticity.
We examine longitudinal dynamics ($\mathbf{u}_{l}$) by taking the divergence of (48) we are examining the dynamics of $\mathbf{u}_{l}$. We find the two coupled equations
$\hbar\left(\frac{\omega^{2}}{\mathcal{E}^{\prime\prime}}-\frac{n-\varrho}{m}\mathbf{k}^{2}\right)\Phi-i\omega\left(n-\varrho+\frac{\mu_{1}}{\mathcal{E}^{\prime\prime}}\right)\gradient\cdot\mathbf{u}_{l}=0\,,\$ (51)
$\begin{split}\left[m\varrho\omega^{2}-\left(K-\frac{\mu_{1}}{\mathcal{E}^{\prime\prime}}\mu_{1}+2G\right)\mathbf{k}^{2}\right]\gradient\cdot\mathbf{u}_{l}&\\ +i\omega\hbar\mathbf{k}^{2}\left(n-\varrho+\frac{\mu_{1}}{\mathcal{E}^{\prime\prime}}\right)&\Phi=0\end{split}$ (52)
in the variables $\mathbf{u}_{l}$ and $\phi$. It can be shown easily that if we are in a superfluid regime the second equation is trivially satisfied, and the first equation finds the Bogoliubov frequency. These equations can be solved for $\omega$ as a function of $\mathbf{k}$ and leads to the following dispersion relation
$\omega^{2}=\frac{A\pm D}{2m\varrho}\,\mathbf{k}^{2}$ (53)
with
$\begin{split}A&=K+2G+\left(n\mathcal{E}^{\prime\prime}+2\mu_{1}\right)\left(n-\varrho\right)\\ D^{2}&=A^{2}-4\varrho(n-\varrho)\left[\mathcal{E}^{\prime\prime}(K+2G)-\mu_{1}^{2}\right]\,.\end{split}$ (54)
We note that, in the case of the homogeneous superfluid ($\varrho=0$), we obtain the superfluid velocity directly from (47)
$\omega^{2}=\frac{n\,U_{k=0}}{m}\mathbf{k}^{2}$ (55)
which is exactly equal to the Bogoliubov velocity. In the opposite limit of $\varrho=n$, the two modes have frequencies
$\omega^{2}=\frac{K+2G}{mn}\mathbf{k}^{2}\quad\text{and zero.}$. (56)
The zero mode corresponds to the collapse of the Bogoliubov mode and the linear dispersing mode has a velocity as expected for a normal solid with elastic moduli $K$ and $G$.
It is interesting to note that, although we expect the Gross-Pitaevskii Lagrangian to struggle with describing a normal solid, the effective phenomenological Lagrangian that emerges actually returns to that of a normal solid in the limit $\varrho\rightarrow n$. Moreover, although all of the modes have a factor of $1/\varrho$ in their velocities which seems to indicate that the velocities should blow up as $\varrho\rightarrow 0$, the velocities appear to be well behaved as $\varrho$ becomes small. This suggests that there is some implicit dependence on $\varrho$ hidden in the values of $B_{iklm}$ etc., which stops the velocities from blowing up. However, we should note that as we appear to see a first-order transition, the value of $\varrho$ never really approaches zero in our numerics, but only jumps directly to zero at which point these equations are no longer valid.
We can also now examine the nature of the normal modes of the system. By writing (51) and (52) in terms of a matrix equation, we can find the eigenvectors of the system for a given $\omega$. These eigenvectors are the normal modes, and are some composite of the fields $\phi$ and $\mathbf{u}_{l}$. We can then subsequently define an eigenvector angle $\theta$ which describes the relative contribution of the fields $\phi$ and $\mathbf{u}_{l}$ to the normal mode. For example, in the superfluid the eigenvector angle is $\theta=0$ as the normal mode is purely $\phi$, i.e. $\mathbf{e}_{n}\equiv\mathbf{e}_{\phi}$. A supersolid with normal mode $\mathbf{e}_{n}=\left(0.5,0.5\right)$ would have $\theta=\pi/4$ etc. The normal mode composition is illustrated in Figure 5.
An interesting point of note is that the slow longitudinal mode (which is expected to die off in the limit of a solid) is a composite of both $\phi$ and $\mathbf{u}_{l}$, but is mostly $\phi$ at all points in the phase diagram (i.e. $\theta$ is small). More importantly, it appears there is a critical point at which the slow longitudinal mode is ’saturated’, that is to say that there is a global maximum in the mixing of the slow longitudinal mode (but note that the mode is still mostly $\phi$). This feature appears to be absent in the fast longitudinal mode. However, the fast longitudinal mode also exhibits interesting behaviour as the mixing angle does not tend to $\pi/2$ or $23\pi/$ 2 as one may initially expect (as we would expect the mode to tend to a purely elastic phonon mode in the limit of a solid), but instead tends to approximately $87\pi/$ 8. This is a curious feature of the fast longitudinal mode, and we are currently investigating the origin of this behaviour.
The above holds for any system described by a Gross-Pitaevskii equation; we can now use these results in the specific case of a soft-core interaction. We find a ground state for the system numerically and by applying the procedure detailed above we can determine the velocities of all modes that exist. Importantly we note that our Lagrangian is tuned across the transition, i.e. a single theory describes both the superfluid and supersolid regime and continuously varies across a first order phase transition. Importantly, symmetry arguments by Son [12] dictate the functional form of any Lagrangian belonging to a supersolid and since our Lagrangian is a subset of those allowed by symmetry, we argue that we completely capture the phenomenological behaviour of the supersolid as described by a Hartree-Fock approximation.
Figure 5: Top: excitation velocities (in units of $\hbar/\sqrt{n}ma^{2}$) for all available sound modes of a system governed by a Heaviside interaction of strength $\Lambda$ with a phase transition at $\Lambda\sim 38$, as calculated via homogenisation. Note that in the superfluid phase the bulk mode is the Bogoliubov mode, hence the sudden appearance of two extra modes at the superfluid-supersolid transtion. Bottom: the eigenvector angle $\theta$ for the available longitudinal modes.
VIII Bogoliubov Dispersion
An alternative technique to finding excitations of the system is to directly apply linearised oscillations in the number density and the phase. Namely, since $\psi=\sqrt{\rho}e^{i\phi}$, we can apply the Bogoliubov fluctuations $\rho\rightarrow\rho_{0}+\delta\rho$ and $\phi$. We expect to recover all the Goldstone modes accessible to the system, and expect the quantitative properties to be the same as those predicted in the elastic theory. More precisely, due to the broken $U(1)$ phase and the broken continuous translational symmetry, we expect to find three Goldstone modes, two of which are longitudinal and one of which is a shear mode. Since this is a completely independent technique to recovering the Goldstone modes, it serves as an important check of our results.
Whilst both techniques access the speeds of sound, the elastic technique also provides elastic moduli and fractional inertia, whereas the bogoliubov technique provides a full band structure. In this way, the two techniques complement each other, as they verify common ground but also provide information that is inaccesible to the other.
By following a similar procedure to that outlined in [14], we are able to obtain eigenvalue equations which give the excitation frequencies $\omega$ of the Bloch waves in the system. These Bloch waves have a band structure, and the lowest band structure at the $\Gamma$ point corresponds to the Goldstone modes of the system whose propagation velocities were independently derived in [14]. The band structure is fully described by the equations
$\begin{split}&\left[-\omega^{2}+\left(\hat{T}^{\mathbf{k}}+\hat{U}^{\mathbf{k}}\right)\hat{T}^{\mathbf{k}}\right](\Phi)^{\mathbf{k}}=0\\ &\left[-\omega^{2}+\hat{T}^{\mathbf{k}}\left(\hat{T}^{\mathbf{k}}+\hat{U}^{\mathbf{k}}\right)\right](\delta\psi)^{\mathbf{k}}=0\end{split}$ (57)
with $\hat{T}$ and $\hat{U}$ defined as
$\begin{split}\hat{T}^{\mathbf{k}}_{\mathbf{G},\mathbf{G}^{\prime}}=&\left[\frac{1}{2}\left(\mathbf{G}+\mathbf{k}\right)^{2}-\mu\right]\delta_{\mathbf{G},\mathbf{G}^{\prime}}+U_{\mathbf{G}-\mathbf{G}^{\prime}}\rho_{\mathbf{G}-\mathbf{G}^{\prime}}\\ \hat{U}^{\mathbf{k}}_{\mathbf{G},\mathbf{G}^{\prime}}=&2\sum_{\mathbf{G}^{\prime\prime}}\psi_{\mathbf{G}-\mathbf{G}^{\prime\prime}}U_{\mathbf{k}+\mathbf{G}^{\prime\prime}}\psi_{\mathbf{G}^{\prime\prime}-\mathbf{G}^{\prime}}\end{split}$ (58)
with $U_{\mathbf{q}}$, $\psi_{\mathbf{q}}$ and $\rho_{\mathbf{q}}$ the fourier transforms of the interaction potential, the superfluid order parameter and the density respectively. We expect that the band structure obtained from solving these equations should find all the requisite Goldstone modes and that they should have the same velocities as predicted via homogenisation.
Figure 6: Band structure for a supersolid at $\Lambda=50$. The three Goldstone modes are highlighted as a fast longitudinal, slow longitudinal and shear mode.
Upon recovery of the band structure, one finds exactly $(D+1)$ Goldstone modes as expected. There is a band crossing at the $K$ point of the fast longitudinal and shear modes, characteristic of a triangular lattice (n.b. this is not actually a crossing but a band touching, as shown by the $M\rightarrow\Gamma$ band structure).
For any particular given $\Lambda$, both the Bogoliubov and elastic techniques predict all three speeds of sound (if supersolid, only one speed if superfluid), and agree on the speeds with excellent precision. Subsequently, by deriving the excitations of the system in two independent ways and obtaining the same results to excellent numerical precision, we can reliably say we have obtained the correct dynamics.
IX Results and Discussion
We can see that we have now verified the technique of homogenisation as a means of recovering the low-lying excitations of the system. This was achieved via means of two completely independent techniques, the results of which both agree to excellent accuracy. An advantage of homogenisation over the Bogoliubov technique is that it also provides the elastic constants of the system, which cannot be obtained via Bogoliubov. Moreover, it provides the fractional inertia of the system, which is a quantity that is of great physical interest. A clear disadvantage however, is that homogenisation does not provide the full spectrum of excitations, only the long-wavelength limit. We stress that this is not an issue, as the Bogoliubov technique is completely separate from homogenisation, and so by utilising both techniques one can recover the full band structure from Bogoliubov and supplement it with additional information in the long-wavelength limit from homogenisation.
The effective Lagrangian that is the end result of homogenisation elucidates interesting dynamics. By considering that the flow excitations of the system can be coupled to the strain excitations, we can see that the system is capable of strain-induced flow. Furthermore, by considering the application of an instantaneous strain to the system, we can see that flow will be induced and contribute to the elastic response of the system. Alternatively, if we apply a strain and keep it fixed for a long time so that the flow dissipates, there will be a different elastic response. In this way we can see that the elastic modulus has a time dependence, and that the system is capable of dissipating energy via the flow of the system. This energy dissipation is likely to be via the two longitudinal modes which are coupled crystal and Bogoliubov phonons.
There is an interesting question of what the compressibility of the system is. When one considers the elastic strain, it is natural to define the compressibility of the system as $\kappa=-1/V\,(\partial V/\partial P)$. Typically in a solid, one would simultaneously be able to say $\kappa=1/n\,(\partial n/\partial P)$, where $n$ is the number density and $\mu$ the chemical potential. However, in our analysis, we can show that these two definitions do not agree, and are off by a considerable margin. These two values are defined as $\kappa=1/B_{xxyy}$ and $\kappa=1/(n\mu_{1})$ in our notation respectively.
The resolution of this seeming dilemma is to revisit the definition of the two compressibilities. The compressibility $\kappa=1/B$ where $B=-V\,(\partial P/\partial V\bigr{|}_{N})$ is the elastic bulk modulus of the system, taken at constant particle number. This is the compressibility of the system as defined via elastic theory, and is the one that we will take as the ’true compressibility’, and consequently $B$ as the ’true bulk modulus’. Importantly, the derivative is taken under constant particle number. As such, to elucidate this a little further, we write
$B_{e}=-V\frac{\partial P}{\partial V}\biggr{|}_{N}$ (59)
where the subscript $e$ denotes that this is the elastic bulk modulus. We can then write the compressibility as $\kappa_{e}=1/B_{e}$. We can subsequently define the thermodynamic bulk modulus (using the typical definition with $n$ but now removing the factor of $V$ which is kept constant) as
$B_{t}=N\frac{\partial P}{\partial N}\biggr{|}_{V}$ (60)
and the thermodynamic compressibility accordingly.
The key consideration is that the $B_{e}\neq B_{t}$ due to the fact that we are keeping different variables constant upon taking derivatives. If we want to connect these two definitions, we need to use the cyclic relations of thermodynamics
$\frac{\partial x}{\partial y}\biggr{|}_{z}=-\frac{\partial x}{\partial z}\biggr{|}_{y}\frac{\partial z}{\partial y}\biggr{|}_{x}\quad.$ (61)
We can then write
$B_{e}=-V\frac{\partial P}{\partial V}\biggr{|}_{N}=V\frac{\partial P}{\partial N}\biggr{|}_{V}\frac{\partial N}{\partial V}\biggr{|}_{P}\equiv B_{t}\ \frac{1}{n}\frac{\partial N}{\partial V}\biggr{|}_{P}$ (62)
from which we can consider the following. In a regular material (solid, gas, liquid) we would expect that the derivative of the particle number with respect to the volume under constant pressure is simply the density of the system, and we would recover the thermodynamic bulk modulus $B_{t}=n\,(\partial P/\partial n\bigr{|}_{V})$ from (62). We can then examine whether this is the case for our system, i.e. is $11/n\,(\partial N/\partial V\bigr{|}_{P})=$ 1?
We can examine this quantity, henceforth denoted as $\mathcal{R}$ and called the ’compressibility ratio’, by choosing an initial volume and particle number $V_{0}$ and $N_{0}$ which will generate a $P_{0}$, adjusting the volume around $V_{0}$ with some $\delta V$, and then varying the particle number $N_{0}+\delta N$ until the system reaches $P_{0}$. We can then numerically calculate $\mathcal{R}$ by finite differencing. If we expect the difference between the two compressibilities to be due to the fact that $\frac{1}{n}\frac{\partial N}{\partial V}\bigr{|}_{P}\neq 1$ then we would expect $\mathcal{R}=\frac{\kappa_{t}}{\kappa_{e}}$, or equivalently
$\mathcal{R}=\frac{B_{e}}{B_{t}}\equiv\frac{K}{n\mu_{1}}\,.$ (63)
Upon performing the numerical determination of $\mathcal{R}$, we find that it is within good agreement of the ratio of the two compressibilities as determined via homogenisation. The percentage error between the homogenisation and numerical $\mathcal{R}$ is typically between $3\%$ and $5\%$. We believe this is an artifact of the numerical determination of $\mathcal{R}$, and that as the algorithm and precision of simulations for determining $\mathcal{R}$ is improved, the agreement will improve accordingly.
Figure 7: Compressibility ratio $\mathcal{R}$ for a system governed by a soft-core interaction of strength $\Lambda$. $\mathcal{R}$ is obtained using the relation (63), where the values are calculated using homogenisation. The expected value of $\mathcal{R}=1$ is obtained in the superfluid phase, but is starkly different in the supersolid regime.
While we have shown that the compressibility ratio obtained via homogenisation is in good agreement with the numerical determination, we do not yet fully understand the nature of this anomalous result. In our search across the literature we have not found another example of this phenomenon, and we are currently investigating the origin of this effect. A possible explanation for the result is considering a pressure which can be written like $P(n,G(n,V))$, where $G$ is the ordering wavevector. One can then subsequently show that the difference between bulk moduli can be written as
$B_{e}=B_{t}-V\frac{\partial P}{\partial G}\biggr{|}_{n}\frac{\partial G}{\partial V}\biggr{|}_{n}$ (64)
where it is clear that the last term is the anomalous term.
This term vanishes in ordinary materials in various ways. For gases and liquids, there is no ordering vector, so this term never existed to begin with. A regular solid has well-defined atomic positions. So, if the density is specified, the ordering wavevector is also specified. Upon changing the volume, there is no additional change in the ordering wavevector beyond that caused by the change in density. Supersolids are unique in that the ordering vector $G$ experiences a commensuration effect. For instance, in a one-dimensional system, it needs to be an integer multiple of $2\pi/L$ with $L$ being the length of the system. If we stretch the system by $\delta L$ much less than the lattice spacing, then the fractional change in the ordering wavevector is $\delta G/G=-\delta L/L$ so that $L\partial G/\partial L=-G$. We believe this is a feature unique to supersolids, due to their ability to exist at a wide range of densities, each of which has their own ordering vector, whereas regular solids only have one density at which they can exist (at some given pressure) and therefore only have one ordering vector. This is (to the best of our knowledge) a new phenomena which could be of great interest to the characterisation of supersolids.
Some currently open questions are as follows. Is this effect unique to the supersolid phase, or is it a more general phenomena? What kind of thermal transport is occurring in the phonon modes of a supersolid? Does the notion of the Landau criterion for superfluidity still hold in the supersolid phase, and if so, how does it change now that we have a band structure? These questions may potentially have avenues to solution via the techniques developed in this paper.
The successful development of homogenisation techniques for the study of supersolids is a significant step forward in the study of these systems. We have shown that the homogenisation technique is able to accurately capture the behaviour of the supersolid phase through comparison with well established Bogoliubov techniques. We were able to obtain previously unknown elastic properties of these materials as well as an effective Lagrangian which can be used to study the dynamics of the system. In contrast with typical phases of matter which have $\mathcal{R}=1$, we found that the compressibility ratio of the supersolid phase is $\mathcal{R}>1$. This is a new phenomena which we predict will arise in supersolids, and of which the origin is currently unknown.
Appendix A Expansion calculation
We start by expanding the non-interacting part of $\mathcal{L}$ using the tools developed. We can express without expansion
$\left(\nabla^{\prime}\rho(\mathbf{r}^{\prime})\right)^{2}=\left[\left(\partial_{i}+\left(u_{ik}+u_{il}u_{lk}\right)\partial_{k}\right)\left(\rho_{0}(\mathbf{r})+\tilde{\rho}(\mathbf{r})\right)\right]^{2}$ (65)
where we subsequently drop the coordinates as it is clear everything is in the unprimed basis $\mathbf{r}$. We can re-express the pre-factor
$\frac{1}{\rho(\mathbf{r}^{\prime})}=\frac{1}{\rho_{0}+\tilde{\rho}}=\frac{1}{\rho_{0}}-\frac{\tilde{\rho}}{\rho_{0}^{2}}+\frac{\tilde{\rho}^{2}}{\rho_{0}^{3}}+\mathop{}\!\mathcal{O}(\tilde{\rho}^{3})$ (66)
to make it more algebraically pliable. This allows us to write
$\int_{\Omega^{\prime}}\frac{\hbar^{2}}{2m}$ $\frac{(\nabla^{\prime}\rho(\mathbf{r}^{\prime}))^{2}}{4\rho(\mathbf{r}^{\prime})}\mathop{}\!\mathrm{d}\mathbf{r}^{\prime}=\int_{\Omega}\Big{\{}\frac{\hbar^{2}}{8m}\left(\frac{1}{\rho_{0}}-\frac{\tilde{\rho}}{\rho_{0}^{2}}+\frac{\tilde{\rho}^{2}}{\rho_{0}^{3}}\right)\left(1+\epsilon_{ll}+M_{iklm}u_{ik}u_{lm}\right)\times$ (67)
$\left[(\nabla\rho_{0})^{2}+2(\gradient\rho_{0}\cdot\gradient\tilde{\rho})-2\epsilon_{ik}\partial_{i}\rho_{0}\partial_{k}\rho_{0}+2\chi_{ik}\partial_{i}\rho_{0}\partial_{k}\rho_{0}-4\epsilon_{ik}\partial_{i}\rho_{0}\partial_{k}\tilde{\rho}+(\nabla\tilde{\rho})^{2}+2\Delta_{ik}\partial_{i}\rho_{0}\partial_{k}\rho_{0}\right]\Big{\}}\mathop{}\!\mathrm{d}\mathbf{r}$
which we can manipulate further. By isolating first and second order (in terms of the strain tensor) parts of the non-interacting Lagrangian, and using the following relation (obtained via integration by parts)
$\int_{\Omega}\frac{\gradient\tilde{\rho}\cdot\gradient\rho_{0}}{\rho_{0}}\mathop{}\!\mathrm{d}\mathbf{r}=\int_{\Omega}\tilde{\rho}\frac{\left(\nabla\rho_{0}\right)^{2}}{\rho_{0}^{2}}-\tilde{\rho}\frac{\nabla^{2}\rho_{0}}{\rho_{0}}\mathop{}\!\mathrm{d}\mathbf{r}$ (68)
we are able to find the first-order component
$\frac{\hbar^{2}}{4m}\left[\left(\frac{(\nabla\rho_{0})^{2}}{2\rho_{0}^{2}}-\frac{\nabla^{2}\rho_{0}}{\rho_{0}}\right)\tilde{\rho}+\epsilon_{ik}\left(\frac{(\nabla\rho_{0})^{2}}{2\rho_{0}}\delta_{ik}-\frac{\partial_{i}\rho_{0}\partial_{k}\rho_{0}}{\rho_{0}}\right)\right]\$ (69)
and the second-order component
$\frac{\hbar^{2}}{4m}\left[\left(\Delta_{ik}+\chi_{ik}\right)\frac{\partial_{i}\rho_{0}\partial_{k}\rho_{0}}{\rho_{0}}\right]+$ (70)
$\frac{\hbar^{2}}{4m}\frac{1}{2}\left[-4\epsilon_{ik}\frac{\partial_{i}\rho_{0}\partial_{k}\tilde{\rho}}{\rho_{0}}\right.+\frac{(\nabla\tilde{\rho})^{2}}{\rho_{0}}+2\frac{\gradient\tilde{\rho}\cdot\gradient\rho_{0}}{\rho_{0}}\epsilon_{ll}-2\epsilon_{ik}\frac{\partial_{i}\rho_{0}\partial_{k}\rho_{0}}{\rho_{0}}\epsilon_{ll}$
$\qquad\quad+\frac{(\nabla\rho_{0})^{2}}{\rho_{0}}\,M_{iklm}u_{ik}u_{lm}-2\frac{\gradient\tilde{\rho}\gradient\rho_{0}}{\rho_{0}^{2}}\tilde{\rho}\left.+2\epsilon_{ik}\frac{\partial_{i}\rho_{0}\partial_{k}\rho_{0}}{\rho_{0}^{2}}\tilde{\rho}-\frac{(\nabla\rho_{0})^{2}}{\rho_{0}^{2}}\tilde{\rho}\epsilon_{ll}+\frac{(\nabla\rho_{0})^{2}}{\rho_{0}^{3}}\tilde{\rho}^{2}\right]$
respectively. We now turn to considering the interaction section of the Lagrangian.
By first noting that the interaction potential is centrosymmetric (and indeed we expect most effective potentials to be isotropic), we can write the interaction as a function of radial distance. We can express the change in the metric length of distance by simply applying the deformation definition to |$\absolutevalue{\Delta\mathbf{r}^{\prime}}^{2}$ and obtain the result
$\left(\Delta r^{\prime}\right)^{2}=\left(\delta_{ik}+2\epsilon_{ik}+2\Delta_{ik}\right)\left(\Delta r\right)_{i}\left(\Delta r\right)_{k}$ (71)
which is exactly equivalent to that of Landau & Lifshitz [28]. The tensor in the above equation is sometimes referred to as the finite strain tensor.
Recalling the derivation of (19) we define the functions
$U(\Delta\mathbf{r}^{\prime})=U(\Delta\mathbf{r})+\left(\epsilon_{ik}+\Delta_{ik}\right)f_{ik}(\Delta\mathbf{r})+\epsilon_{ik}\epsilon_{lm}W_{iklm}(\Delta\mathbf{r})$ (72)
with
$f_{ik}(\mathbf{r})\equiv\frac{r_{i}r_{k}}{|\mathbf{r}|}\frac{\partial U(|\mathbf{r}|)}{\partial|\mathbf{r}|}\quad;\quad W_{iklm}(\mathbf{r})\equiv\frac{r_{i}r_{k}r_{l}r_{m}}{2|\mathbf{r}|^{2}}\left(\frac{\partial^{2}U}{\partial|\mathbf{r^{2}}|}-\frac{1}{|\mathbf{r}|}\frac{\partial U}{\partial|\mathbf{r}|}\right)\;.$ (73)
This can be used to express the interaction component of $\mathcal{L}$ by combining it with both the density mapping and the Jacobian to obtain
$U=\frac{1}{2}\int_{\Omega}\left(\rho_{0}+\tilde{\rho}\right)_{r_{1}}\left[U(\mathbf{r}_{12})+\left(\epsilon_{ik}+\Delta_{ik}\right)f_{ik}(\mathbf{r}_{12})\right.$ $\left.+\epsilon_{ik}\epsilon_{lm}W_{iklm}(\mathbf{r}_{12})\right]\left(\rho_{0}+\tilde{\rho}\right)_{r_{2}}$ (74)
$\left(1+2\epsilon_{ll}+\epsilon_{ll}\epsilon_{kk}+2M_{iklm}u_{ik}u_{lm}\right)\;\mathop{}\!\mathrm{d}\mathbf{r}_{1}\mathop{}\!\mathrm{d}\mathbf{r}_{2}\quad$
where $\mathbf{r}_{12}=\mathbf{r}_{1}-\mathbf{r}_{2}$. We note that, since the convolution is integrated, there is a term arising from $\left(\mathcal{J}_{\mathbf{r}^{\prime}\rightarrow\mathbf{r}}\right)^{2}$.
Once again, this can be expanded and combined with the non-interacting component of $\mathcal{L}$, and we can fully express the first and second order components of the expansion. We make use of the Gross-Pitaevskii ground-state equation (2), as well as the normalisation condition (14) to simplify our expressions, and we can show that
$\mathcal{L}_{1}=\int_{\Omega}\left[\rho_{0}\epsilon_{ll}\left(U*\rho_{0}-\mu_{0}\right)+\frac{1}{2}\epsilon_{ik}(f_{ik}*\rho_{0})\rho_{0}+\frac{\hbar^{2}}{4m}\epsilon_{ik}\left(\frac{\left(\nabla\rho_{0}\right)^{2}}{2\rho_{0}}\delta_{ik}-\frac{\partial_{i}\rho_{0}\partial_{k}\rho_{0}}{\rho_{0}}\right)\right]\mathop{}\!\mathrm{d}\mathbf{r}$ (75)
We can now make a direct connection between the linear component of the expansion and the pressure tensor $P_{ik}$. Recalling that the first order component of the expansion is the stress tensor, and that the stress tensor is the negative of the pressure, we can write
$\begin{split}P_{ik}&=\frac{1}{|\Omega|}\int_{\Omega}\left[\left(\mu_{0}-U*\rho_{0}\right)\delta_{ik}-\frac{1}{2}f_{ik}*\rho_{0}-\frac{\hbar^{2}}{4m}\left(\frac{\left(\nabla\rho_{0}\right)^{2}}{2\rho_{0}^{2}}\delta_{ik}-\frac{\partial_{i}\rho_{0}\partial_{k}\rho_{0}}{\rho_{0}^{2}}\right)\right]\rho_{0}\mathop{}\!\mathrm{d}\mathbf{r}\\ &=\frac{1}{|\Omega|}\int_{\Omega}\left[\frac{\hbar^{2}}{4m}\left(\frac{\partial_{i}\rho_{0}\partial_{k}\rho_{0}}{\rho_{0}}-\nabla^{2}\rho_{0}\,\delta_{ik}\right)-\frac{1}{2}\rho_{0}(f_{ik}*\rho_{0})\right]\mathop{}\!\mathrm{d}\mathbf{r}\\ \end{split}$ (76)
[…]
It is important to note that since we have allowed $\tilde{\rho}$ to have both first and second order components in $u_{ik}$, then even single powers of $\tilde{\rho}$ must be taken into consideration for the $\mathop{}\!\mathcal{O}(u^{2})$ components. The above analysis only collected the components of $\tilde{\rho}$ which were of $\mathop{}\!\mathcal{O}(u)$, but does not remove the singular powers of $\tilde{\rho}$ from the Lagrangian. It is necessary to calculate the functional form of $\tilde{\rho}$ as there will be terms like $\int_{\Omega}\tilde{\rho}^{2}\mathop{}\!\mathrm{d}\mathbf{r}$ that will contribute to the perturbative Lagrangian, and as such we therefore need to solve the equations of motion for $\tilde{\rho}$. It is not good enough to simply know the normalisation condition as we did for the $\mathop{}\!\mathcal{O}(u)$ component.
The first step in solving for $\tilde{\rho}$ is to collect all the terms in which it appears, regardless of first or second order splitting. We take all the terms that contain $\tilde{\rho}$ and call them $\mathcal{L}_{\tilde{\rho}}$, finding
$-\mathcal{L}_{\tilde{\rho}}=$ $\frac{1}{2}\frac{\hbar^{2}}{4m}\left[-4\epsilon_{ik}\frac{\partial_{i}\rho_{0}\partial_{k}\tilde{\rho}}{\rho_{0}}+\frac{(\nabla\tilde{\rho})^{2}}{\rho_{0}}+2\frac{\gradient\tilde{\rho}\cdot\gradient\rho_{0}}{\rho_{0}}\epsilon_{ll}-2\frac{\gradient\tilde{\rho}\cdot\gradient\rho_{0}}{\rho_{0}^{2}}\tilde{\rho}+2\epsilon_{ik}\frac{\partial_{i}\rho_{0}\partial_{k}\rho_{0}}{\rho_{0}^{2}}\tilde{\rho}-\frac{(\nabla\rho_{0})^{2}}{\rho_{0}^{2}}\tilde{\rho}\,\epsilon_{ll}+\frac{(\nabla\rho_{0})^{2}}{\rho_{0}^{3}}\tilde{\rho}^{2}\right]$ (77)
$+\mu_{0}\tilde{\rho}+\frac{1}{2}\left(\left(U*\tilde{\rho}\right)\tilde{\rho}+4\epsilon_{ll}\left(U*\rho_{0}\right)\tilde{\rho}+2\epsilon_{ik}\left(f_{ik}*\rho_{0}\right)\tilde{\rho}\right)$
which we can solve by the standard Euler-Lagrange technique. A key detail now is that the new normalisation condition changes the constraints we apply. Specifically, the equation we need to solve is
$\frac{\partial\mathcal{L}_{\tilde{\rho}}}{\partial\tilde{\rho}}-\partial_{i}\frac{\partial\mathcal{L}_{\tilde{\rho}}}{\partial\left(\partial_{i}\tilde{\rho}\right)}+\left(\mu_{0}+\mu_{1}\epsilon_{ll}+\mu_{2}^{iklm}u_{ik}u_{lm}\right)\left(1+\epsilon_{ll}+M_{iklm}u_{ik}u_{lm}\right)=0$ (78)
which can be shown via the strained normalisation condition (14).
A straightforward calculation follows of differentiating the $\tilde{\rho}$ dependent Lagrangian and collating all terms. We make use of the zeroth-order equation of motion to simplify our expression and we find
$\begin{split}&\frac{\hbar^{2}}{4m}\gradient\cdot\left(\frac{\gradient\tilde{\rho}}{\rho_{0}}\right)-U*\tilde{\rho}\\[5.0pt] +&\frac{\hbar^{2}}{4m}\left(\frac{(\nabla\rho_{0})^{2}}{\rho_{0}^{3}}-\frac{\nabla^{2}\rho_{0}}{\rho_{0}^{2}}\right)\tilde{\rho}\end{split}\text{\Large{\hbox{ = }}}\begin{split}&-\left(\mu_{2}^{iklm}+\mu_{1}^{ik}\delta_{lm}+\mu_{0}M_{iklm}\right)u_{ik}u_{lm}+\epsilon_{ik}\left[-\mu_{1}^{ik}+\frac{\hbar^{2}}{4m}\left(2\frac{\partial_{ik}\rho_{0}}{\rho_{0}}-\frac{\partial_{i}\rho_{0}\partial_{k}\rho_{0}}{\rho_{0}^{2}}\right)\right.\\[5.0pt] &\left.+\int_{\Omega}\left(f_{ik}(\mathbf{r}-\mathbf{r}^{\prime})+\delta_{ik}U(\mathbf{r}-\mathbf{r}^{\prime})\right)\rho_{0}(\mathbf{r}^{\prime})\mathop{}\!\mathrm{d}\mathbf{r}^{\prime}\right]\end{split}$ (79)
which contains both $\mathop{}\!\mathcal{O}(u^{2})$ and $\mathop{}\!\mathcal{O}(u)$ terms. We can now begin to separate $\tilde{\rho}$ into first and second order components via (13), and consider solving the equation component by component. That is to say that we solve the lowest order component first, and then use that solution to solve the next order component, and so on.
The first order equation to solve is given by
$\begin{split}&\frac{\hbar^{2}}{4m}\gradient\cdot\left(\frac{\gradient\left(\rho_{1}^{ik}\epsilon_{ik}\right)}{\rho_{0}}\right)-U*\left(\rho_{1}^{ik}\epsilon_{ik}\right)\\ &+\frac{\hbar^{2}}{4m}\left(\frac{(\nabla\rho_{0})^{2}}{\rho_{0}^{3}}-\frac{\nabla^{2}\rho_{0}}{\rho_{0}^{2}}\right)\left(\rho_{1}^{ik}\epsilon_{ik}\right)\end{split}\text{\Large{\hbox{ = }}}\begin{split}\epsilon_{ik}&\left[-\mu_{1}^{ik}+\frac{\hbar^{2}}{4m}\left(2\frac{\partial_{ik}\rho_{0}}{\rho_{0}}-\frac{\partial_{i}\rho_{0}\partial_{k}\rho_{0}}{\rho_{0}^{2}}\right)\right.\\[5.0pt] &\left.+\int_{\Omega}\left(f_{ik}(\mathbf{r}-\mathbf{r}^{\prime})+\delta_{ik}U(\mathbf{r}-\mathbf{r}^{\prime})\right)\rho_{0}(\mathbf{r}^{\prime})\mathop{}\!\mathrm{d}\mathbf{r}^{\prime}\right]\end{split}$ (80)
which we can solve numerically. We factor out $\epsilon_{ik}$ and solve the set of equations where we have one for each index (in 2D we have 4 equations to solve, in 3D there are 9). As this is an integrodifferential equation we can solve the equation by writing the LHS as a matrix operator on an unraveled vector and applying an iterative matrix solver algorithm which converges on some pre-specified tolerance. We can introduce notation and write the first and second order equations of motion as
$L\left(\rho_{1}^{ik}\right)=F_{1}\left(\rho_{0},\mu_{1}^{ik}\right)\,,\qquad L\left(\rho_{2}^{iklm}\right)=F_{2}\left(\mu_{0},\mu_{1}^{ik},\mu_{2}^{iklm}\right)$ (81)
for brevity. It is important to note that we do not actually need to solve the equation for $\rho_{2}^{iklm}$ as it only appears in the Lagrangian when multiplied by a constant, thus we can directly use the normalisation condition (14) and short circuit the task of solving it. We also do not need to find $\mu_{2}^{iklm}$ for a slightly different reason, which is that it appears in the Lagrangian to too high order and so is always discarded anyway. Thus our only task now is to determine the first order correction to the chemical potential $\mu_{1}^{ik}$ and the first order solution to $\tilde{\rho}$ which is $\rho_{1}^{ik}$.
We first state $\rho_{1}^{ik}\equiv\rho_{1}^{ik}(\mathbf{r},\mu_{1}^{ik})$, i.e. it is now a function of the chemical potential. We specifically choose the functional form $\rho_{1}^{ik}=\hat{\rho}_{1}^{ik}+\mu_{1}^{ik}\rho_{1}^{\prime}$, in a kind of ’homogeneous + inhomogeneous’ fashion. Namely, we demand that
$L\left(\hat{\rho}_{1}^{ik}\right)=\frac{\hbar^{2}}{4m}\left(2\frac{\partial_{ik}\rho_{0}}{\rho_{0}}-\frac{\partial_{i}\rho_{0}\partial_{k}\rho_{0}}{\rho_{0}^{2}}\right)+\int_{\Omega}\left[f_{ik}(\mathbf{r}-\mathbf{r}^{\prime})+\delta_{ik}U(\mathbf{r}-\mathbf{r}^{\prime})\right]\rho_{0}(\mathbf{r}^{\prime})\mathop{}\!\mathrm{d}\mathbf{r}^{\prime}$ (82)
i.e. $\hat{\rho}_{1}^{ik}$ satisfies the equation of motion without any chemical potential considerations. Then we solve the equation
$L\left(\mu_{1}^{ik}\rho_{1}^{\prime}\right)=-\mu_{1}^{ik}\,\rightarrow\,L\left(\rho_{1}^{\prime}\right)=-1$ (83)
simultaneously, which we can do numerically as it is the same operator but just a different RHS. We can state this in the notation of (81)
$L\left(\hat{\rho}_{1}^{ik}\right)=F_{1}\left(\rho_{0},\mu_{1}^{ik}=0\right)\quad,\quad L\left(\rho_{1}^{\prime}\right)=F_{1}\left(\rho_{0}=0,\mu_{1}^{ik}=1\right)$ (84)
which we can readily substitute into the ’homogeneous’ and ’inhomogeneous’ solutions to see that
$L\left(\hat{\rho}_{1}^{ik}+\mu_{1}^{ik}\rho_{1}^{\prime}\right)=F_{1}\left(\rho_{0},\mu_{1}^{ik}\right)$ (85)
which is the ‘general’ solution we were looking for. The final task is to now actually calculate $\mu_{1}^{ik}$ which we can do by using the normalisation condition(14). Since we know what the first-order condition is, we can rearrange to find that
$\mu_{1}^{ik}=-\int_{\Omega}(\rho_{0}\delta_{ik}+\hat{\rho}_{1}^{ik})\mathop{}\!\mathrm{d}\mathbf{r}\Bigg{/}\int_{\Omega}\rho_{1}^{\prime}\mathop{}\!\mathrm{d}\mathbf{r}$ (86)
which is the final ingredient.
We have now fully solved for $\tilde{\rho}$ (and discarded the parts that we don’t need), so we can re-express $\mathcal{L}_{\tilde{\rho}}$ using integration by parts. The calculation is relatively straightforward and makes use of the base EoM (2), allowing one to find
$-\mathcal{L}_{\tilde{\rho}}=\frac{1}{2}\left(2\mu_{0}\rho_{2}^{iklm}u_{ik}u_{lm}+\rho_{1}^{ik}\left[\frac{\hbar^{2}}{4m}\left(2\frac{\partial_{ik}\rho_{0}}{\rho_{0}}-\frac{\partial_{i}\rho_{0}\partial_{k}\rho_{0}}{\rho_{0}^{2}}\right)+\left(f_{ik}+\delta_{ik}U\right)*\rho_{0}+2\mu_{0}\delta_{ik}+\mu_{1}^{ik}\right]\,\epsilon_{ik}\epsilon_{lm}\right)+\mathop{}\!\mathcal{O}(\epsilon^{3})$ (87)
which contains all the second-order contributions from $\tilde{\rho}$.
We can combine this with all the second-order contributions that do not contain $\tilde{\rho}$, and obtain a long expression
$\begin{split}-\mathcal{L}_{2}=\frac{1}{2}&\left[2\left(U*\rho_{0}\right)\rho_{0}M_{iklm}u_{ik}u_{lm}+\left(\left(U*\rho_{0}\right)\rho_{0}\delta_{ik}\delta_{lm}+2\left(f_{ik}*\rho_{0}\right)\rho_{0}\delta_{lm}+\left(W_{iklm}*\rho_{0}\right)\rho_{0}\right)\,\epsilon_{ik}\epsilon_{lm}\right.\\ &\left.+\left(f_{ik}*\rho_{0}\right)\rho_{0}\Delta_{ik}+2\frac{\hbar^{2}}{4m}\frac{\partial_{i}\rho_{0}\partial_{k}\rho_{0}}{\rho_{0}}\left(\Delta_{ik}+\chi_{ik}\right)+\frac{\hbar^{2}}{4m}\left(-2\frac{\partial_{i}\rho_{0}\partial_{k}\rho_{0}}{\rho_{0}}\right)\delta_{lm}\epsilon_{ik}\epsilon_{lm}\right.\\ &\left.+\frac{\hbar^{2}}{4m}\frac{\left(\nabla\rho_{0}\right)^{2}}{\rho_{0}}M_{iklm}u_{ik}u_{lm}+2\mu_{0}\rho_{0}\left(\delta_{ik}\delta_{lm}-M_{iklm}\right)u_{ik}u_{lm}\right.\\ &\left.+\rho_{1}^{lm}\left[\frac{\hbar^{2}}{4m}\left(2\frac{\partial_{ik}\rho_{0}}{\rho_{0}}-\frac{\partial_{i}\rho_{0}\partial_{k}\rho_{0}}{\rho_{0}^{2}}\right)+\left(f_{ik}+\delta_{ik}U\right)*\rho_{0}+2\mu_{0}\delta_{ik}+\mu_{1}^{ik}\right]\,\epsilon_{ik}\epsilon_{lm}\right]\,.\end{split}$ (88)
which will be used in the calculation of the elastic constants.
To summarise, we have expanded our Lagrangian up to second order in the strain tensor and collected all the terms. We’ve solved for $\tilde{\rho}$ such that the new state is a ground state with the strained normalisation condition, and subsequently re-expressed the second-order Lagrangian with that solution in hand.
Appendix B Connecting Cauchy elastic tensor to expansion
This section closely follows the calculation done by Bavaud et al [29]. To begin with, we define an analytic expansion in the free energy of our system around a deformation $x_{\alpha}^{\prime}=\left(\delta_{\alpha\beta}-u_{\alpha\beta}\right)x_{\beta}$ with a mapping $\Omega\mapsto\Omega^{\prime}$
$F(\Omega^{\prime})=F(\Omega)-\absolutevalue{\Omega}\pi_{\alpha\beta}(\Omega)u_{\alpha\beta}+\absolutevalue{\Omega}\frac{1}{2}A_{\alpha\beta\gamma\delta}(\Omega)u_{\alpha\beta}u_{\gamma\delta}$ (89)
where
$\pi_{\alpha\beta}(\Omega)=\left.\frac{1}{\absolutevalue{\Omega}}\frac{\partial F(\Omega^{\prime})}{\partial\left(-u_{\alpha\beta}\right)}\right|_{u=0}=\left.-\frac{1}{\absolutevalue{\Omega}}\frac{\partial F(\Omega^{\prime})}{\partial u_{\alpha\beta}}\right|_{u=0}$ (90)
and similarly
$A_{\alpha\beta\gamma\delta}(\Omega)=\left.\frac{1}{\absolutevalue{\Omega}}\frac{\partial^{2}F(\Omega^{\prime})}{\partial u_{\alpha\beta}u_{\gamma\delta}}\right|_{u=0}\quad.$ (91)
To clarify, we map between spaces $\Omega\mapsto\Omega^{\prime}$ through the transformation $\mathbf{x}\mapsto\mathbf{x^{\prime}}$.
It is important to note that the tensor $A_{\alpha\beta\gamma\delta}$ is not the elastic tensor. It is the expansion of the free energy due to the deformation, but the elastic tensor strictly connects the stress tensor to the strain. We are searching for Hooke’s Law which maps strain in a deformed point to strain in an undeformed point but now with the notation
$\pi_{ik}(\Omega^{\prime})=\pi_{ik}(\Omega)-B_{iklm}(\Omega)u_{lm}\quad.$ (92)
An important additional feature of the elastic tensor $B_{iklm}$ is that, for a sufficiently symmetric system, it has certain symmetry properties as described by Landau & Lifshitz [28], whereas the expansion tensor $A_{iklm}$ does not necessarily have these symmetries.
We first define an additional two deformations, $u^{\prime}$ and $u^{\prime\prime}$ such that
$x_{\alpha}^{\prime\prime}=\left(\delta_{\alpha\beta}-u_{\alpha\beta}^{\prime}\right)x_{\beta}^{\prime}\quad;\quad x_{\alpha}^{\prime\prime}=\left(\delta_{\alpha\beta}-u_{\alpha\beta}^{\prime\prime}\right)x_{\beta}$ (93)
To clarify, we have an initial space $\Omega$ which we can deform with the transformation $\mathbf{u}$ into the space $\Omega^{\prime}$. We can then deform from $\Omega^{\prime}\mapsto\Omega^{\prime\prime}$ with the deformation $\mathbf{u}^{\prime}$, and we can also deform to the $\Omega^{\prime\prime}$ space from the original space like $\Omega\mapsto\Omega^{\prime\prime}$ with the deformation $\mathbf{u}^{\prime\prime}$. We can combine the deformations to find the relation
$u_{\alpha\beta}^{\prime\prime}=u_{\alpha\beta}+u_{\alpha\beta}^{\prime}-u_{\alpha\sigma}^{\prime}u_{\sigma\beta}\quad.$ (94)
We then use the above relation to expand the stress tensor that connects the $\Omega^{\prime}$ and $\Omega^{\prime\prime}$ space. Examining the expansion of the free energy from the $\Omega^{\prime}$ space to the $\Omega^{\prime\prime}$ space, we find
$F(\Omega^{\prime\prime})=F(\Omega^{\prime})-\absolutevalue{\Omega^{\prime}}\pi_{\alpha\beta}(\Omega^{\prime})u_{\alpha\beta}^{\prime}+\frac{\absolutevalue{\Omega^{\prime}}}{2}A_{\alpha\beta\gamma\delta}(\Omega^{\prime})u_{\alpha\beta}^{\prime}u_{\gamma\delta}^{\prime}\,.$ (95)
Expanding $\pi_{\alpha\beta}(\Omega^{\prime})$ and using (90) and (94), we find
$\begin{split}\pi_{\alpha\beta}\left(\Omega^{\prime}\right)=&\left.-\frac{1}{\absolutevalue{\Omega^{\prime}}}\frac{\partial F(\Omega^{\prime\prime})}{\partial u_{\alpha\beta}^{\prime}}\right|_{u^{\prime}=0}\\ =&\left.-\frac{1}{\absolutevalue{\Omega^{\prime}}}\frac{\partial F(\Omega^{\prime\prime})}{\partial u_{\gamma\epsilon}^{\prime\prime}}\frac{\partial u_{\gamma\epsilon}^{\prime\prime}}{\partial u_{\alpha\beta}^{\prime}}\right|_{u^{\prime}=0}\\ =&\left.-\frac{1}{\absolutevalue{\Omega^{\prime}}}\frac{\partial F(\Omega^{\prime\prime})}{\partial u_{\gamma\epsilon}^{\prime\prime}}\left(\delta_{\gamma\alpha}\delta_{\beta\epsilon}-\delta_{\gamma\alpha}\delta_{\beta\sigma}u_{\sigma\epsilon}\right)\right|_{u^{\prime}=0}\\ =&\left.-\frac{1}{\absolutevalue{\Omega^{\prime}}}\left(\delta_{\beta\epsilon}-u_{\beta\epsilon}\right)\frac{\partial F(\Omega^{\prime\prime})}{\partial u_{\alpha\epsilon}^{\prime\prime}}\right|_{u^{\prime}=0}\,.\end{split}$ (96)
We can now turn our attention to the derivative in the above expression. Namely, since the derivative actually does not depend on $u^{\prime}$ explicitly (recall that we can map from $\Omega\mapsto\Omega^{\prime\prime}$ through $u^{\prime\prime}$ without needing $u^{\prime}$ at all), we can remove the evalutation at $u^{\prime}=0$. As such, the derivative is now not a number, but a function instead! This function describes the energy response of the mapping $\Omega\mapsto\Omega^{\prime\prime}$ through $u^{\prime\prime}$ for any arbitrary $\Omega^{\prime\prime}$ and $u^{\prime\prime}$. That means it must be equivalent to the function which maps from $\Omega\mapsto\Omega^{\prime}$ through $u$. Thus we can write
$\left.\frac{\partial F(\Omega^{\prime\prime})}{\partial u_{\alpha\epsilon}^{\prime\prime}}\right|_{u^{\prime}=0}=\frac{\partial F(\Omega^{\prime})}{\partial u_{\alpha\epsilon}}$ (97)
and the final result from (96) becomes
$\pi_{\alpha\beta}\left(\Omega^{\prime}\right)=-\frac{1}{\absolutevalue{\Omega^{\prime}}}\left(\delta_{\beta\epsilon}-u_{\beta\epsilon}\right)\frac{\partial F(\Omega^{\prime})}{\partial u_{\alpha\epsilon}}\quad.$ (98)
Since the derivative is no longer evaluated at $u=0$, we now substitute in (89) to find
$\frac{\partial F(\Omega^{\prime})}{\partial u_{\alpha\epsilon}}=-\absolutevalue{\Omega}\pi_{\alpha\epsilon}(\Omega)+\absolutevalue{\Omega}A_{\alpha\epsilon\gamma\delta}(\Omega)u_{\gamma\delta}\quad.$ (99)
We now also expand the $1/\absolutevalue{\Omega^{\prime}}$ | term to find
[…]
where the indices $x$ and $y$ are tensor indices of the same nature as $\alpha$, $\beta$ etc. It is important to note that we are discarding terms of $\mathop{}\!\mathcal{O}(u^{2})$ or higher as we are trying to derive Hooke’s law, i.e. $F=kx$, and we do not want anharmonic terms.
We can now combine the results of (99) and (100) into (98) (discarding $\mathop{}\!\mathcal{O}(u_{ik}^{2})$ terms) to find
$\pi_{\alpha\beta}(\Omega^{\prime})=\pi_{\alpha\beta}-\left(A_{\alpha\beta\gamma\delta}+\delta_{\gamma\beta}\pi_{\alpha\delta}-\delta_{\gamma\delta}\pi_{\alpha\beta}\right)u_{\gamma\delta}$ (101)
where everything on the RHS is a function of $\Omega$. This is the expression of Hooke’s Law as described in (92)! We can therefore find that
$B_{iklm}=A_{iklm}+\pi_{im}\delta_{kl}-\pi_{ik}\delta_{lm}$ (102)
with now the final task being to relate $\pi_{ik}$ to the pressure tensor $P_{ik}$. Recalling that the stress tensor is just the negative of the pressure, we can write the final expression as
$B_{iklm}=A_{iklm}+P_{ik}\delta_{lm}-P_{im}\delta_{kl}$ (103)
which is the Cauchy elastic tensor. For the case of $C_{3}$ symmetry, we have the relation [28]:
$B_{xxxx}=B_{xxyy}+2B_{xyxy}\,.$ (104)
Appendix C Tensor differentiation
We have shown in previous sections how to expand a Lagrangian under a strain tensor. Once we have that expansion, how do we recover the homogenised tensors as first and second differentives: $\pi_{ik}=\partial F/\partial u_{ik}$ and $A_{iklm}=\partial^{2}F/\partial u_{ik}\partial u_{lm}$?
We will now write out explicitly how expressions involving the tensors listed in (9) differentiate. Consider first $F=F_{ik}u_{ik}$ with $F_{ik}=F_{ki}$, its first derivative is: $\partial F/\partial u{ik}=F_{ik}$. Similarly, the first derivative of $F=F_{ik}\epsilon_{ik}$ is $\partial F/\partial u_{ik}=-F_{ik}$.
Let us now consider second derivatives. Suppose there is a term in the Lagrangian of the form $F=F_{iklm}u_{ik}u_{lm}$. We have the result $\partial^{2}F/\partial u_{ik}\partial u_{lm}=2F_{iklm}$. Consider now $F=F_{ik}\chi_{ik}$ where $\chi_{ik}=\left(u_{li}u_{kl}+u_{lk}u_{il}\right)/2$. Since $\chi_{ik}$ is symmetric under $i\leftrightarrow k$, we can impose that $F_{ik}$ is also symmetric under the same exchange.
$\frac{\partial^{2}\chi_{ik}}{\partial u_{cd}\partial u_{ab}}$ $=\frac{1}{2}\frac{\partial}{\partial u_{cd}}\left(\delta_{la}\delta_{ib}u_{kl}+\delta_{la}\delta_{kb}u_{il}+\delta_{ka}\delta_{lb}u_{li}+\delta_{ia}\delta_{lb}u_{lk}\right)=\frac{1}{2}\left(\delta_{ib}\delta_{kc}\delta_{ad}+\delta_{kb}\delta_{ci}\delta_{ad}+\delta_{ka}\delta_{bc}\delta_{id}+\delta_{ia}\delta_{bc}\delta_{kd}\right)\,.$ (105)
This allows us to obtain
$\frac{\partial^{2}F}{\partial u_{ik}\partial u_{lm}}=\frac{1}{2}\left(F_{bc}\delta_{ad}+F_{cb}\delta_{ad}+F_{da}\delta_{bc}+F_{ad}\delta_{bc}\right)=F_{kl}\delta_{im}+F_{im}\delta_{kl}\,.$ (106)
using the symmetry of $F_{ik}$ under exchange of indices. A similar calculation can be done for other tensor contractions involving the tensors in (9). This is listed in Table 1.
Table 1:
$F$ $F_{iklm}\epsilon_{ik}\epsilon_{lm}$ $F_{ik}\Delta_{ik}$ $F_{ik}\omega_{ik}$ $F_{ik}\chi_{ik}$
${\frac{\partial^{2}F}{\partial u_{ik}\partial u_{lm}}}$ $2F_{iklm}$ $F_{il}\delta_{km}$ $F_{km}\delta_{il}$ $F_{kl}\delta_{im}+F_{im}\delta_{kl}$
It is important to briefly discuss the symmetry of each of these expressions. For terms that look like $F_{iklm}\epsilon_{ik}\epsilon_{lm}$, these already have all the symmetry properties we would want from our elastic tensors. However, we can immediately see that the contribution from $F_{ik}\Delta_{ik}$, $F_{ik}\omega_{ik}$ and $F_{ik}\chi_{ik}$ do not have the requisite symmetries. We instead need to create these symmetries from combinations of these tensors. This cannot be achieved a priori by adding in whatever we like. However, if we expect elastic theory to hold, then we would expect any expansion in any strain tensor (including ours) to lead us eventually back to symmetric expressions. This is a good test to check whether the expansion calculation has been done correctly.
Appendix D Numerical Deformations
In our numerics, we use a spatial discretisation consisting of a triangular lattice as a grid so that the numerical scheme is consistent with the symmetry of the supersolid phase under rotations. Therefore, each “pixel” is a parallelogram and the simulation cell is also a parallelogram with basis vectors $\mathbf{a}$ and $\mathbf{b}$ parallel to $(\sqrt{3}/2,1/2)$ and $(0,1)$ in Cartesian coordinates. This appendix provides some algebraic details on how deformations $\mathbf{V}$ of the simulation cell can be used to extract the elastic tensors numerically using the energy change $F_{2}\equiv Pu_{ll}+\frac{1}{2}A_{iklm}u_{ik}u_{lm}$ up to second order in the strain.
Let us write all positions and displacements in the basis of the triangular grid: $\mathbf{r}=r_{a}\mathbf{a}+r_{b}\mathbf{b}$ and $\mathbf{u}=u_{a}\mathbf{a}+u_{b}\mathbf{b}$. The strain tensor in this basis
$\mathbf{V}=\begin{pmatrix}\partial u_{a}/\partial r_{a}&\partial u_{a}/\partial r_{b}\\ \partial u_{b}/\partial r_{a}&\partial u_{b}/\partial r_{b}\end{pmatrix}$
is related to the strain tensor in the Cartesian basis: $(\mathbf{U})_{ij}\equiv u_{ij}$ by:
$\mathbf{U}=\mathbf{T}^{-1}\mathbf{V}\mathbf{T}\,,\qquad\mbox{with}\qquad\mathbf{T}=\begin{pmatrix}\frac{2}{\sqrt{3}}&0\\ -\frac{1}{\sqrt{3}}&1\end{pmatrix}\,.$ (107)
In our numerics, we apply strains diagonal in the triangular-grid basis by deforming the simulation cell in the directions of $\mathbf{a}$ and $\mathbf{b}$. Using the above, we can calculate the energy changes we should expect to see. There are three deformations that give linearly independent results.
$\mathbf{V}=\begin{pmatrix}\epsilon&0\\ 0&0\end{pmatrix}\!:F_{2}$ $=P\epsilon+\left(A_{xxxx}+\frac{A_{xyxy}}{3}\right)\frac{\epsilon^{2}}{2}$
$\mathbf{V}=\begin{pmatrix}0&0\\ 0&\epsilon\end{pmatrix}\!:F_{2}$ $=P\epsilon+\left(A_{yyyy}+\frac{A_{xyxy}}{3}\right)\frac{\epsilon^{2}}{2}$ (108)
$\mathbf{V}=\begin{pmatrix}\epsilon&0\\ 0&\epsilon\end{pmatrix}\!:F_{2}$ $=2P\epsilon+\left(\frac{A_{xxxx}+A_{yyyy}}{2}+A_{xxyy}\right)\epsilon^{2}\,.$
For our system with $C_{3}$ symmetry, there are only two independent elastic constants. Using (103) and (104), we see that $A_{xxxx}=A_{xxyy}+P+2A_{xyxy}$. Therefore, we have three equations to solve for the elastic constants and the pressure for a given $\epsilon\ll 1$.
We have calculated these quantities at several values of $\epsilon$ to check for the self-consistency of this analysis. This gives a measure of the numerical errors of this technique.
A good question to ask is whether or not we are allowed to a priori assume the system is symmetric, and whether we are biasing towards a symmetric result. This is a valid question, but really what we are doing is seeking to compare the results we obtain with the analytical theory derived. Importantly, the theory assumes nothing about symmetry, so if the numerics agree with the theory, then this assumption is justified.